### v2

# Local Asymptotic Normality for Mixed Fractional Brownian Motion under High-Frequency Observation

Abstract
In this paper we will consider the LAN property for both the Hurst parameter $H>3/4$ and the variance of the fractional Brownian motion plus an independent standard Brownian motion (called mixed fractional Brownian motion) with high-frequency observation. We will first remove the $H$-score linear term and orthogonalize the remainder through two non-diagonal transformations, then we can construct the CLT for the quadratic form base on $\|\cdot\|_{\mathrm{op}}/\|\cdot\|_{F}\to 0$. At last we obtain a diagonal Gaussian LAN expansion with an explicit information matrix. Beyond the case of $H>3/4$, we also present that the $\|\cdot\|_{\mathrm{op}}/\|\cdot\|_{F}\to 0$ method is also useful for the case of $H<3/4$ and the proof will be concise compared with the Whittle translation method. We consider that this method can be applied to this type of problem, including the fractional Ornstein-Uhlenbeck model and mixed fractiona O-U process.
1 Introduction
We consider the mixed fractional Brownian motion (mfBm)
$Y_{t}=\sigma B_{t}^{H}+W_{t},\qquad t\geq 0,$ (1)
where $B^{H}$ is a standard fractional Brownian motion with Hurst index $H\in(3/4,1)$, $W$ is an independent standard Brownian motion, and the parameter $\theta=(\sigma,H)$ is unknown. We observe discrete high-frequency increments on a growing time horizon:
$\Delta_{n}=n^{-\alpha},\quad\alpha\in(0,1),\qquad\mathcal{T}_{n}:=n\Delta_{n}\longrightarrow\infty,$
$X_{n,i}:=Y_{i\Delta_{n}}-Y_{(i-1)\Delta_{n}}=\sigma\Delta_{n}^{H}G_{n,i}+\Delta_{n}^{1/2}Z_{n,i},\qquad i=1,\dots,n,$ (2)
where $(G_{n,i})$ is fractional Gaussian noise (fGn) and $(Z_{n,i})$ is standard Gaussian white noise, independent of $(G_{n,i})$. In this paper we will construct the LAN property for the unknown parameter $\theta$ under this high-frequency observation.
We now recall Le Cam’s notion of local asymptotic normality (LAN), which provides a precise second-order approximation of the log-likelihood ratio under local perturbations and yields sharp minimax lower bounds.
Definition 1 (LAN).
Let $\bigl(\mathbf{P}_{\theta}^{n}\bigr)_{\theta\in\Theta}$ be a sequence of statistical experiments indexed by $\theta\in\Theta\subset\mathbb{R}^{k}$. We say it is locally asymptotically normal at $\theta_{0}\in\Theta$ if there exist non-singular $k\times k$ matrices $\phi(n)=\phi(n,\theta_{0})$ with $\|\phi(n)\|\to 0$, random vectors $Z_{n,\theta_{0}}\in\mathbb{R}^{k}$, and a symmetric nonnegative definite matrix $I(\theta_{0})$ such that for every fixed $u\in\mathbb{R}^{k}$,
$\log\frac{d\mathbf{P}^{n}_{\theta_{0}+\phi(n)u}}{d\mathbf{P}^{n}_{\theta_{0}}}(X^{n})=u^{\top}Z_{n,\theta_{0}}-\frac{1}{2}u^{\top}I(\theta_{0})u+o_{\mathbb{P}_{\theta_{0}}}(1),\qquad Z_{n,\theta_{0}}\Rightarrow\mathcal{N}\bigl(0,I(\theta_{0})\bigr).$
Denote by $W_{2,k}$ the class of loss functions $\ell:\mathbb{R}^{k}\to[0,\infty)$ that are symmetric, subconvex and satisfy
$\lim_{\|u\|\to\infty}e^{-a\|u\|^{2}}\,\ell(u)=0,\qquad\forall a>0.$
The following is the classical local asymptotic minimax lower bound for LAN families.
Theorem 1.
Assume $\bigl(\mathbf{P}_{\theta}^{n}\bigr)_{\theta\in\Theta}$ is LAN at $\theta_{0}$ with normalization matrices $\phi(n)$ and information $I(\theta_{0})$. Then for any sequence of estimators $\hat{\theta}_{n}$, any $\ell\in W_{2,k}$, and any $\delta>0$,
$\liminf_{n\to\infty}\;\sup_{\|\theta-\theta_{0}\|<\delta}\mathbf{E}_{\theta}\!\left[\ell\!\left(\phi(n)^{-1}\bigl(\hat{\theta}_{n}-\theta\bigr)\right)\right]\ \geq\ \int_{\mathbb{R}^{k}}\ell(x)\,\varphi_{I(\theta_{0})^{-1}}(x)\,dx,$
where $\varphi_{I(\theta_{0})^{-1}}$ denotes the centered Gaussian density with covariance $I(\theta_{0})^{-1}$ (restricted to the range of $I(\theta_{0})$ when $I(\theta_{0})$ is singular).
In a continuous-noise setting Kleptsyna and Chigansky [6] have obtained LAN-type statements under continuous observation.A natural question is whether one can simply “transfer” continuous-observation LAN results by replacing the continuous time horizon with the discrete one $\mathcal{T}_{n}$.
However, in the present infill scheme (2), the observation is inherently discrete and contaminated by the Brownian component at the increment level: the small-scale behavior is dominated by the $\Delta_{n}^{1/2}$-term, while the long-memory component enters through a vanishing regularization factor (in our notation later $\gamma_{n}=\sigma^{2}\Delta_{n}^{2H-1}\to 0$). This mismatch changes both the effective Fisher information scaling and the form of the local reparametrization.
In [10], the author has considered the same model when the long-memory singularity is mild (notably $H<3/4$), Whittle-type approximations can be used to translate Toeplitz likelihoods into spectral integrals and to establish LAN in a comparatively direct way. However, the regime $H>3/4$ corresponds to a stronger low-frequency singularity: the relevant trace functionals are driven by the near-zero behavior of the spectral density of fGn $f_{H}$ and its $H$-derivative and require refined trace asymptotics for triangular-array (i.e. $n$-dependent) regularization.
The most delicate phenomenon in the regime $H>3/4$ is that the (properly normalized) score vector exhibits an additional asymptotic co-linearity (see the Appendix) which is not present in the pure fGn model [9]. This degeneracy arises from a specific interplay between the vanishing signal-to-noise ratio and the severity of the spectral singularity. Unlike pure fGn, where self-similarity preserves the signal structure across scales, the mfBm signal is asymptotically buried under white noise because the fractional variance $\Delta_{n}^{2H}$ decays faster than the white noise variance $\Delta_{n}$ whenever $H>1/2$. However, noise dominance alone does not imply degeneracy; in the intermediate regime $1/2<H<3/4$, where the noise also dominates, the LAN property can be established without projection because the spectral density is square-integrable (the memory parameter p=2H1<1/2p=2H-1<1/2). In stark contrast, when $H>3/4$, we have $p>1/2$, rendering the squared spectral density non-integrable at the origin. This strong singularity concentrates the Fisher information so heavily at zero frequency that the spectral “shape” variations (governed by $H$) become asymptotically indistinguishable from the “energy” scaling (governed by $\sigma$). Consequently, the standard score vectors collapse into a rank-1 subspace, necessitating our projection step to recover the orthogonal information.
The rest of the paper will be organized as follows: in Section 2 we will introduce the main results of the LAN property. Section 3 to 5 are contributed to the proof of our theorem. Then we will the simulated results in Section 6. At last in the Appendix, some explicit formula in the Fisher information will be introduced and we will also explain why we should take two non-diagonal rate matrix for this case. Also in the Appendix we extend our $\|\cdot\|_{\mathrm{op}}/\|\cdot\|_{F}\to 0$ method to the case $H<3/4$, this method will be more concise than the Whittle translation method presented in [10].
2 Main Results
For any integrable (real-valued) symbol $u$ on $\pi,\pi]$, we define $T_{n}(u)$ be the $n\times n$ Toeplitz matrix with entries
$(T_{n}(u))_{jk}:=\hat{u}_{j-k},\qquad\hat{u}_{k}:=\frac{1}{2\pi}\int_{-\pi}^{\pi}u(\lambda)e^{-ik\lambda}\,d\lambda,\quad k\in\mathbb{Z}.$
In particular, $T_{n}(H)=T_{n}(f_{H})$, where $f_{H}$ is the spectral density of the standard fGn under the Fourier inversion
$\rho_{H}(k)=\frac{1}{2\pi}\int_{-\pi}^{\pi}e^{ik\lambda}f_{H}(\lambda)\,d\lambda.$
Also as we know $f_{H}(\lambda)$ has the low-frequency asymptotic
$f_{H}(\lambda)\sim c_{H}|\lambda|^{-(2H-1)}\qquad(\lambda\to 0),$ (3)
for some constant $c_{H}>0$ depending only on $H$. We define the signal-to-noise ratio parameter
$\gamma_{n}=\sigma^{2}\Delta_{n}^{2H-1},$ (4)
then our observation $X^{n}=(X_{n,1},\dots,X_{n,n})^{\top}$ is a centered Gaussian vector with covariance
$V_{n}(\theta)=\Delta_{n}\big(I_{n}+\gamma_{n}T_{n}(H)\big),\qquad\theta=(\sigma,H),$ (5)
where $I_{n}$ is the $n\times n$ identity matrix and $T_{n}(H)$ is the Toeplitz covariance matrix of the standard fractional Gaussian noise (fGn) with Hurst index $H\in(1/2,1)$. We consider the increasing-horizon high-frequency sampling scheme
$\Delta_{n}=n^{-\alpha},\qquad 0<\alpha<1,\qquad\mathcal{T}_{n}:=n\Delta_{n}=n^{1-\alpha}\to\infty.$ (6)
2.1 Exact score functions
Let $\ell_{n}(\theta)$ be the Gaussian log-likelihood of $X_{n}\sim\mathcal{N}(0,V_{n}(\theta))$:
$\ell_{n}(\theta)=-\frac{n}{2}\ln(2\pi)-\frac{1}{2}\ln\det V_{n}(\theta)-\frac{1}{2}X_{n}^{\top}V_{n}(\theta)^{-1}X_{n},$
then
Proposition 1.
The exact score components are centered quadratic forms
$S_{\sigma,n}:=\partial_{\sigma}\ell_{n}$ $=\frac{\gamma_{n}}{\sigma}\Big(X_{n}^{\top}\Psi_{\sigma,n}X_{n}-\mathbb{E}[X_{n}^{\top}\Psi_{\sigma,n}X_{n}]\Big),$ (7)
$S_{H,n}:=\partial_{H}\ell_{n}$ $=\frac{\gamma_{n}}{2}\Big(X_{n}^{\top}\Psi_{H,n}X_{n}-\mathbb{E}[X_{n}^{\top}\Psi_{H,n}X_{n}]\Big),$ (8)
where
$\Psi_{\sigma,n}$ $=(I_{n}+\gamma_{n}T_{n}(H))^{-1}\,T_{n}(H)\,(I_{n}+\gamma_{n}T_{n}(H))^{-1},$
$\Psi_{H,n}$ $=(I_{n}+\gamma_{n}T_{n}(H))^{-1}\,\big(2\ln(\Delta_{n})\,T_{n}(H)+\dot{T}_{n}(H)\big)\,(I_{n}+\gamma_{n}T_{n}(H))^{-1},$
and $\dot{T}_{n}(H):=\partial_{H}T_{n}(H)$.
Proof.
This is the standard Gaussian score identity:
$\partial_{\theta}\ell_{n}=-\frac{1}{2}\mathrm{tr}(V_{n}^{-1}\partial_{\theta}V_{n})+\frac{1}{2}X_{n}^{\top}V_{n}^{-1}(\partial_{\theta}V_{n})V_{n}^{-1}X_{n},$
and $\mathbb{E}[X_{n}^{\top}AX_{n}]=\mathrm{tr}(AV_{n})$. Using $V_{n}=\Delta_{n}(I_{n}+\gamma_{n}T_{n})$ together with $\partial_{\sigma}V_{n}=\frac{2\gamma_{n}\Delta_{n}}{\sigma}T_{n}$ and $\partial_{H}V_{n}=\Delta_{n}\gamma_{n}(2\ln(\Delta_{n})T_{n}+\dot{T}_{n})$ yields the stated sandwich forms. ∎
Theorem 2.
Let the rate matrix $M_{n}=M_{n}^{(2)}M_{n}^{(1)}$ defined in (50), the sequence of statistical experiments generated by the observations $X_{n}$ is Locally Asymptotically Normal at $\theta=(\sigma,H)$ for $\sigma>0,\,3/4<H<1$. Specially for any fixed $h\in\mathbb{R}^{2}$, the log-likelihood ratio satisfies:
$\log\frac{d\mathbb{P}^{n}_{\theta+hr_{n}^{-1}}}{d\mathbb{P}_{\theta}^{n}}=h^{\top}\Xi_{n}-\frac{1}{2}h^{\top}I^{\perp}h+o_{\mathbb{P}}(1),$
where
$r_{n}:=\sqrt{\mathcal{T}_{n}}\,(M_{n}^{-1})^{\!\top},\qquad\text{equivalently}\qquad r_{n}^{-\,\top}=\frac{1}{\sqrt{\mathcal{T}_{n}}}\,M_{n}.$
The random vector $\Xi_{n}$ converges in law to $\mathcal{N}(0,I^{\perp})$ and
$I^{\perp}=\begin{pmatrix}\frac{\sigma^{2}}{\pi}J_{0}(H,\sigma)&0\\[5.69054pt] 0&\frac{\sigma^{4}}{4\pi}J_{\perp}(H,\sigma)\end{pmatrix},\quad J_{\perp}=J_{2}-\frac{J_{1}^{2}}{J_{0}},$
with p=2H1p=2H-1 and $J_{1},J_{2}$ defined in (29)–(30).
3 Preliminaries: Score Functions
3.1 Some General Properties of fGn and Score Function
As presented in the introduction, we deal with the score function using the standard normal quadratic formula, so first, we will reduce the it to the standard one. We present
Lemma 1.
Let $A_{n}:=I_{n}+\gamma_{n}T_{n}(H)$ and let $Z\sim\mathcal{N}(0,I_{n})$. Then we can write
$X_{n}=\sqrt{\Delta_{n}}\,A_{n}^{1/2}Z.$
The $\sigma$-score admits the exact representation
$S_{\sigma,n}=\frac{\gamma_{n}}{\sigma}\Big(Z^{\top}C_{n}Z-\mathrm{tr}(C_{n})\Big),$ (9)
and the $H$-score satisfies the decomposition
$S_{H,n}=\sigma\ln(\Delta_{n})\,S_{\sigma,n}+R_{H,n},$ (10)
where
$R_{H,n}:=\frac{\gamma_{n}}{2}\Big(Z^{\top}D_{n}Z-\mathrm{tr}(D_{n})\Big),$ (11)
with the matrices
$C_{n}:=A_{n}^{-1/2}T_{n}(H)A_{n}^{-1/2},\qquad D_{n}:=A_{n}^{-1/2}\dot{T}_{n}(H)A_{n}^{-1/2}.$
Proof.
By the exact score representation, the weighting matrix for the $H$-score corresponds to $2\ln(\Delta_{n})\,T_{n}(H)+\dot{T}_{n}(H)$ sandwiched by $A_{n}^{-1/2}$. Hence, the centered quadratic form splits into the sum of the $2\ln(\Delta_{n})$-part and the $\dot{T}_{n}(H)$-part. Using the identity $\partial_{H}\gamma_{n}=2\gamma_{n}\ln(\Delta_{n})$ and the scaling relation $S_{\sigma,n}=\frac{\gamma_{n}}{\sigma}Q_{n}(C_{n})$, we obtain the decomposition (10) with $R_{H,n}$ as defined in (11). ∎
We present in the following lemma our important procedure to obtain the central limit theorem for the standard quadratic formula, in this lemma the assumption $\left\lVert\cdot\right\rVert_{\mathrm{op}}/\left\lVert\cdot\right\rVert_{F}\to 0$ play important role:
Lemma 2.
Let $Z\sim\mathcal{N}(0,I_{n})$ and let $M_{n}$ be a sequence of real symmetric matrices. Set $Q_{n}(M_{n}):=Z^{\top}M_{n}Z-\mathrm{tr}(M_{n})$. If
$\frac{\left\lVert M_{n}\right\rVert_{\mathrm{op}}}{\left\lVert M_{n}\right\rVert_{F}}\longrightarrow 0,$ (12)
where $\left\lVert\cdot\right\rVert_{\mathrm{op}}$ is the operator norm (or spectral norm) and $\left\lVert\cdot\right\rVert_{F}$ is the Hilbert–Schmidt (or Frobenius) norm, then
$\frac{Q_{n}(M_{n})}{\sqrt{2}\,\left\lVert M_{n}\right\rVert_{F}}\xRightarrow[n\to\infty]{}\mathcal{N}(0,1).$ (13)
Proof.
Diagonalize $M_{n}=U_{n}^{\top}\Lambda_{n}U_{n}$ with $U_{n}$ orthogonal and $\Lambda_{n}=\mathrm{diag}(\lambda_{n,1},\dots,\lambda_{n,n})$. Then $U_{n}Z\stackrel{{\scriptstyle d}}{{=}}Z$ and
$Q_{n}(M_{n})=\sum_{i=1}^{n}\lambda_{n,i}(Z_{i}^{2}-1),\qquad\mathrm{Var}(Q_{n}(M_{n}))=2\sum_{i=1}^{n}\lambda_{n,i}^{2}=2\left\lVert M_{n}\right\rVert_{F}^{2}.$
Let $\xi_{n,i}:=\lambda_{n,i}(Z_{i}^{2}-1)/(\sqrt{2}\left\lVert M_{n}\right\rVert_{F})$, so that $\sum_{i}\xi_{n,i}=\frac{Q_{n}(M_{n})}{\sqrt{2}\left\lVert M_{n}\right\rVert_{F}}$ and the $\xi_{n,i}$ are independent and centered. Since $Z_{i}^{2}-1$ has a finite fourth moment, there exists a constant $C$ such that $\mathbb{E}[(Z_{i}^{2}-1)^{4}]\leq C$. Thus, we can verify the Lyapunov condition:
$\sum_{i=1}^{n}\mathbb{E}|\xi_{n,i}|^{4}\leq C\cdot\frac{\sum_{i}\lambda_{n,i}^{4}}{\left(\sum_{i}\lambda_{n,i}^{2}\right)^{2}}\leq C\left(\frac{\max_{i}|\lambda_{n,i}|}{\sqrt{\sum_{i}\lambda_{n,i}^{2}}}\right)^{2}=C\left(\frac{\left\lVert M_{n}\right\rVert_{\mathrm{op}}}{\left\lVert M_{n}\right\rVert_{F}}\right)^{2}\to 0,$
where we used $\sum_{i}\lambda_{n,i}^{4}\leq(\max_{i}\lambda_{n,i}^{2})\sum_{i}\lambda_{n,i}^{2}$. Thus the Lindeberg–Feller CLT yields (13). A general treatment of CLTs for quadratic forms is in de Jong [3]. ∎
Because the Toeplitz covariance matrix can generated by the spectral density, we need to find the properties of the spectral density of the fractional Gaussnian noise and its derivative with respect to $H$. First is the Fisher–Hartwig representation:
Lemma 3.
Fix $H\in(3/4,1)$ and set $p:=2H-1\in(1/2,1)$. Let $f_{H}$ be the spectral density of the standard fractional Gaussian noise (fGn), i.e. the stationary Gaussian sequence with autocovariance $\rho_{H}(k)=\frac{1}{2}\big(|k+1|^{2H}-2|k|^{2H}+|k-1|^{2H}\big)$. Then $f_{H}$ admits the Fisher–Hartwig representation
$f_{H}(\lambda)=|2\sin(\lambda/2)|^{-p}\,\ell_{H}(\lambda),\qquad\lambda\in[-\pi,\pi],$ (14)
where $\ell_{H}$ is even, $\ell_{H}\in C^{2}([-\pi,\pi])$, and $0<\inf_{\lambda\in[-\pi,\pi]}\ell_{H}(\lambda)\leq\sup_{\lambda\in[-\pi,\pi]}\ell_{H}(\lambda)<\infty$. Moreover, the derivative in $H$ can be written as
$\dot{f}_{H}(\lambda)=f_{H}(\lambda)\,\bigl(-2\ln|2\sin(\lambda/2)|\bigr)+|2\sin(\lambda/2)|^{-p}\,\dot{\ell}_{H}(\lambda),$ (15)
with $\dot{\ell}_{H}$ even and $\dot{\ell}_{H}\in C^{1}([-\pi,\pi])$.
Proof.
A classical explicit representation of the spectral density of standard fGn is (see, e.g., [7, 4])
$f_{H}(\lambda)=c_{H}\,(1-\cos\lambda)\Big(|\lambda|^{-2H-1}+B(\lambda,H)\Big),\qquad\lambda\in[-\pi,\pi],$ (16)
where $c_{H}>0$ depends only on $H$, and
$B(\lambda,H):=\sum_{j=1}^{\infty}\Big((2\pi j+\lambda)^{-2H-1}+(2\pi j-\lambda)^{-2H-1}\Big).$ (17)
The series in (17) converges uniformly on $\pi,\pi]$, and so do its $\lambda$-derivatives up to order $2$ and the $H$-derivative up to order $1$ (because $(2\pi j\pm\lambda)^{-2H-1}$ is $O(j^{-2H-1})$ and its $H$-derivative is $O(j^{-2H-1}\log j)$, both summable for $H>0$). Hence $B(\cdot,H)\in C^{2}([-\pi,\pi])$ and $\partial_{H}B(\cdot,H)\in C^{1}([-\pi,\pi])$, and both are even in $\lambda$.
Using $1-\cos\lambda=2\sin^{2}(\lambda/2)$ and p=2H1p=2H-1, define
$\ell_{H}(\lambda):=|2\sin(\lambda/2)|^{p}\,f_{H}(\lambda).$
Plugging (16) into this definition yields
$\ell_{H}(\lambda)=c_{H}\,|2\sin(\lambda/2)|^{p+2}\Big(|\lambda|^{-2H-1}+B(\lambda,H)\Big)=c_{H}\,|2\sin(\lambda/2)|^{2H+1}\Big(|\lambda|^{-2H-1}+B(\lambda,H)\Big).$
Write
$|2\sin(\lambda/2)|^{2H+1}|\lambda|^{-2H-1}=\Big(\frac{|2\sin(\lambda/2)|}{|\lambda|}\Big)^{2H+1}.$
Since $\sin x/x$ is analytic near 0, the function $\lambda\mapsto\big(|2\sin(\lambda/2)|/|\lambda|\big)^{2H+1}$ extends to a $C^{\infty}$ even function on $\pi,\pi]$ (with value $1$ at $\lambda=0$). The remaining term $|2\sin(\lambda/2)|^{2H+1}B(\lambda,H)$ is $C^{2}$ on $\pi,\pi]$ because $|2\sin(\lambda/2)|^{2H+1}$ is $C^{\infty}$ and $B(\cdot,H)\in C^{2}$. Therefore $\ell_{H}\in C^{2}([-\pi,\pi])$ and is even. Moreover $\ell_{H}(\lambda)>0$ for all $\lambda$ (since $c_{H}>0$, $1-\cos\lambda\geq 0$, and $|\lambda|^{-2H-1}+B(\lambda,H)>0$), hence by continuity on the compact interval $\pi,\pi]$ we have $0<\inf\ell_{H}\leq\sup\ell_{H}<\infty$. This proves (14).
Finally, differentiating (14) with respect to $H$ and using $\dot{p}=2$ gives
$\dot{f}_{H}(\lambda)=\partial_{H}\!\big(|2\sin(\lambda/2)|^{-p}\big)\,\ell_{H}(\lambda)+|2\sin(\lambda/2)|^{-p}\,\dot{\ell}_{H}(\lambda)=|2\sin(\lambda/2)|^{-p}\Big(-2\ln|2\sin(\lambda/2)|\,\ell_{H}(\lambda)+\dot{\ell}_{H}(\lambda)\Big),$
which is exactly (15). The regularity $\dot{\ell}_{H}\in C^{1}$ and evenness follow from the above uniform convergence arguments for $\partial_{H}B$ and the smoothness of $\sin(\cdot)$. ∎
Now, we will propose a generalized Szego trace approximation for triangular arrays.
Proposition 2.
Let $a_{n}(\lambda):=1+\gamma_{n}f_{H}(\lambda)$, $A_{n}=T_{n}(a_{n})$ which is the Toeplitz matrix driven by the function $a_{n}(\lambda)$, for $n\Delta_{n}\to\infty$ and $b,c\in\{f_{H},\dot{f}_{H}\}$, let us define $k(f_{H}):=0$ and $k(\dot{f}_{H}):=1$ which counts the number of logarithmic Fisher–Hartwig factors (present in $\dot{f}_{H}$ but not in $f_{H}(\lambda)$)hence it governs the possible $\log n$ loss in the remainder. Then
$\mathrm{tr}\!\big(T_{n}(b)\,A_{n}^{-1}\,T_{n}(c)\,A_{n}^{-1}\big)=\frac{n}{2\pi}\int_{-\pi}^{\pi}\frac{b(\lambda)c(\lambda)}{a_{n}(\lambda)^{2}}\,d\lambda+R_{n}(b,c),$ (18)
where the remainder satisfies
$R_{n}(b,c)=O\!\big(n\,\log^{\,k(b)+k(c)}n\big),\qquad n\to\infty.$ (19)
Proof.
The symbols $f_{H}$ and $\dot{f}_{H}$ have a single Fisher–Hartwig singularity at the origin; in particular, by Lemma 3, $f_{H}(\lambda)=|2\sin(\lambda/2)|^{-p}\ell_{H}(\lambda)$ with $\ell_{H}\in C^{2}$ bounded away from 0 and $\infty$, and $\dot{f}_{H}$ decomposes into a logarithmic Fisher–Hartwig part $f_{H}(\lambda)\,(-2\log|2\sin(\lambda/2)|)$ plus a smooth Fisher–Hartwig part $|2\sin(\lambda/2)|^{-p}\dot{\ell}_{H}(\lambda)$. The regularization $a_{n}^{-1}=(1+\gamma_{n}f_{H})^{-1}$ is $n$-dependent and introduces a cutoff at frequency scale $\lambda\asymp\gamma_{n}^{1/p}$.
The trace approximation (18)–(19) is a standard consequence of generalized Szegő-type theorems for long-memory Toeplitz matrices under triangular-array regularization (infill asymptotics). One can derive it, for instance, from the main trace approximation results in Dahlhaus [7], which treat ratios of the form $b/(1+\gamma_{n}f)$ and allow logarithmic Fisher–Hartwig factors. An eigenvalue-sum representation and the corresponding remainder control in the quadratic-variation regime are also discussed in Istas–Lang [8]. ∎
We verify the hypotheses for fGn
Lemma 4.
For $b,c\in\{f_{H},\dot{f}_{H}\}$ the integrands $bc/a_{n}^{2}$ in (18) belong to $L^{1}([-\pi,\pi])$, and near $\lambda=0$ one has the uniform bounds
$\frac{|b(\lambda)c(\lambda)|}{a_{n}(\lambda)^{2}}\leq C\,\frac{f_{H}(\lambda)^{2}\bigl(1+|\log|\lambda||\bigr)^{k(b)+k(c)}}{\bigl(1+\gamma_{n}f_{H}(\lambda)\bigr)^{2}},\qquad|\lambda|\leq 1,$ (20)
for a constant $C$ independent of $n$.
Proof.
The integrability and the bound (20) follow from Lemma 3. Indeed, on $|\lambda|\leq 1$ we have $|2\sin(\lambda/2)|\asymp|\lambda|$ and $\ell_{H},\dot{\ell}_{H}$ are bounded, so $f_{H}(\lambda)\asymp|\lambda|^{-p}$ and
$|\dot{f}_{H}(\lambda)|\leq f_{H}(\lambda)\,\bigl(2|\log|\lambda||+C\bigr)+C\,|\lambda|^{-p}\leq C\,f_{H}(\lambda)\bigl(1+|\log|\lambda||\bigr).$
This yields (20). Since $p=2H-1\in(1/2,1)$, the right-hand side of (20) is integrable on [1,1][-1,1] for each $n$: the factor $(1+\gamma_{n}f_{H})^{-2}$ regularizes the $|\lambda|^{-2p}$ singularity and effectively cuts it off at the scale $|\lambda|\asymp\gamma_{n}^{1/p}$. On $\{1<|\lambda|\leq\pi\}$ all symbols are bounded and $a_{n}(\lambda)\geq 1$, hence $bc/a_{n}^{2}\in L^{1}([-\pi,\pi])$. ∎
In the following lemma, we will give the trace approximation defined in Lemma 1. This approximation will be always used in the structure of $\left\lVert\cdot\right\rVert_{\mathrm{op}}/\left\lVert\cdot\right\rVert_{F}\to 0$.
Lemma 5.
Let $\Delta_{n}=n^{-\alpha}$ with $\alpha\in(0,1)$, and define $\gamma_{n}:=\sigma^{2}\Delta_{n}^{p}$ with p=2H1p=2H-1. Then the following trace approximations hold:
$\mathrm{tr}(C_{n}^{2})$ $=\frac{n}{2\pi}\int_{-\pi}^{\pi}g_{n}(\lambda)^{2}\,d\lambda+o\!\Bigl(n\Delta_{n}^{1-2p}\Bigr),$ (21)
$\mathrm{tr}(C_{n}D_{n})$ $=\frac{n}{2\pi}\int_{-\pi}^{\pi}g_{n}(\lambda)h_{n}(\lambda)\,d\lambda+o\!\Bigl(n\Delta_{n}^{1-2p}\ln(1/\Delta_{n})\Bigr),$ (22)
$\mathrm{tr}(D_{n}^{2})$ $=\frac{n}{2\pi}\int_{-\pi}^{\pi}h_{n}(\lambda)^{2}\,d\lambda+o\!\Bigl(n\Delta_{n}^{1-2p}\ln^{2}(1/\Delta_{n})\Bigr),$ (23)
where
$g_{n}(\lambda)=\frac{f_{H}(\lambda)}{a_{n}(\lambda)},\qquad h_{n}(\lambda)=\frac{\dot{f}_{H}(\lambda)}{a_{n}(\lambda)}.$
Proof.
Give a simple symbol $T_{n}(H)=T_{n}:=T_{n}(f_{H})$ and $\dot{T}_{n}(H)=\dot{T}_{n}:=T_{n}(\dot{f}_{H})$. Set $a_{n}(\lambda):=1+\gamma_{n}f_{H}(\lambda)$ so that $A_{n}=I_{n}+\gamma_{n}T_{n}=T_{n}(a_{n}(\lambda))$.
Since $A_{n}=I_{n}+\gamma_{n}T_{n}$ is a polynomial in $T_{n}$, the matrices $A_{n}$ and $T_{n}$ commute, and so do $A_{n}^{-1}$ and $T_{n}$. Using the cyclicity of the trace and the identities $C_{n}=A_{n}^{-1/2}T_{n}A_{n}^{-1/2}$ and $D_{n}=A_{n}^{-1/2}\dot{T}_{n}A_{n}^{-1/2}$, we obtain
$\mathrm{tr}(C_{n}^{2})$ $=\mathrm{tr}\!\big(A_{n}^{-1}T_{n}A_{n}^{-1}T_{n}\big)=\mathrm{tr}\!\big(T_{n}\,A_{n}^{-1}\,T_{n}\,A_{n}^{-1}\big),$ (24)
$\mathrm{tr}(C_{n}D_{n})$ $=\mathrm{tr}\!\big(A_{n}^{-1}T_{n}A_{n}^{-1}\dot{T}_{n}\big)=\mathrm{tr}\!\big(T_{n}\,A_{n}^{-1}\,\dot{T}_{n}\,A_{n}^{-1}\big),$ (25)
$\mathrm{tr}(D_{n}^{2})$ $=\mathrm{tr}\!\big(A_{n}^{-1}\dot{T}_{n}A_{n}^{-1}\dot{T}_{n}\big)=\mathrm{tr}\!\big(\dot{T}_{n}\,A_{n}^{-1}\,\dot{T}_{n}\,A_{n}^{-1}\big).$ (26)
For (24)–(26) we take Proposition 2 and then obtain
$\mathrm{tr}(C_{n}^{2})$ $=\frac{n}{2\pi}\int_{-\pi}^{\pi}\frac{f_{H}(\lambda)^{2}}{a_{n}(\lambda)^{2}}\,d\lambda+O(n)=\frac{n}{2\pi}\int_{-\pi}^{\pi}g_{n}(\lambda)^{2}\,d\lambda+O(n),$
$\mathrm{tr}(C_{n}D_{n})$ $=\frac{n}{2\pi}\int_{-\pi}^{\pi}\frac{f_{H}(\lambda)\dot{f}_{H}(\lambda)}{a_{n}(\lambda)^{2}}\,d\lambda+O(n\log n)=\frac{n}{2\pi}\int_{-\pi}^{\pi}g_{n}(\lambda)h_{n}(\lambda)\,d\lambda+O(n\log n),$
$\mathrm{tr}(D_{n}^{2})$ $=\frac{n}{2\pi}\int_{-\pi}^{\pi}\frac{\dot{f}_{H}(\lambda)^{2}}{a_{n}(\lambda)^{2}}\,d\lambda+O(n\log^{2}n)=\frac{n}{2\pi}\int_{-\pi}^{\pi}h_{n}(\lambda)^{2}\,d\lambda+O(n\log^{2}n).$
Now when $H>3/4$ we have p=2H1>1/2p=2H-1>1/2, hence 2p1>02p-1>0 and
$\frac{n}{n\Delta_{n}^{1-2p}}=\Delta_{n}^{2p-1}\longrightarrow 0.$
Moreover $\log(1/\Delta_{n})=\alpha\log n$, so for $m\in\{1,2\}$,
$\frac{n\log^{m}n}{n\Delta_{n}^{1-2p}\log^{m}(1/\Delta_{n})}=\Delta_{n}^{2p-1}\cdot\frac{\log^{m}n}{(\alpha\log n)^{m}}\longrightarrow 0$
which yields (21)–(23). ∎
3.2 Central Limit Theorem for $S_{\sigma,n}$
In this subsection and next one we will try to use the $\left\lVert\cdot\right\rVert_{\mathrm{op}}/\left\lVert\cdot\right\rVert_{F}\to 0$ to get the CLT of the $S_{\sigma,n}$ and Remainder for $H$. First let us define the weight function on $\mathbb{R}$:
$w(x):=\left(\frac{c_{H}|x|^{-p}}{1+\sigma^{2}c_{H}|x|^{-p}}\right)^{2}.$ (27)
Since $p>1/2$, one has $w\in L^{1}(\mathbb{R})$ and also $w(x)\ln^{k}|x|\in L^{1}(\mathbb{R})$ for $k=1,2$. Define the integral constants:
$J_{0}(H,\sigma)$ $:=\int_{\mathbb{R}}w(x)\,dx,$ (28)
$J_{1}(H,\sigma)$ $:=\int_{\mathbb{R}}w(x)\Bigl(C_{H}-2\ln|x|\Bigr)\,dx,$ (29)
$J_{2}(H,\sigma)$ $:=\int_{\mathbb{R}}w(x)\Bigl(C_{H}-2\ln|x|\Bigr)^{2}\,dx.$ (30)
All three integrals are finite and $J_{0}>0$.
Remark 1.
These integrals arise from the asymptotics of the trace approximations derived in Lemma 5 and will play a crucial role in the explicit formula for the Fisher Information matrix.
By Lemma 3 and the relation $|2\sin(\lambda/2)|\sim|\lambda|$ as $\lambda\to 0$, the spectral density of standard fGn, $f_{H}(\lambda)$, admits the low-frequency expansions:
$f_{H}(\lambda)=c_{H}|\lambda|^{-p}\bigl(1+r_{H}(\lambda)\bigr),\qquad r_{H}(\lambda)\to 0\ \ (\lambda\to 0),$ (31)
and
$\dot{f}_{H}(\lambda)=f_{H}(\lambda)\Bigl(C_{H}-2\ln|\lambda|+\rho_{H}(\lambda)\Bigr),\qquad\rho_{H}(\lambda)\to 0\ \ (\lambda\to 0),$ (32)
for some constants $c_{H}>0$ and $C_{H}\in\mathbb{R}$ depending on $H$. We now establish the asymptotics for $\mathrm{tr}(C_{n}^{2})$.
Lemma 6.
Under the expansion (31) and the trace approximation (21), we have
$\mathrm{tr}(C_{n}^{2})=\frac{n}{2\pi}\Delta_{n}^{1-2p}\Bigl(J_{0}(H,\sigma)+o(1)\Bigr).$ (33)
Proof.
By (21), it suffices to analyze the integral $I_{0,n}:=\int_{-\pi}^{\pi}g_{n}(\lambda)^{2}\,d\lambda$. Fix a small $\eta\in(0,\pi)$ and split the integral into $I_{0,n}=I_{0,n}^{(0)}+I_{0,n}^{(\infty)}$ with
$I_{0,n}^{(0)}:=\int_{|\lambda|\leq\eta}g_{n}(\lambda)^{2}\,d\lambda,\qquad I_{0,n}^{(\infty)}:=\int_{\eta<|\lambda|\leq\pi}g_{n}(\lambda)^{2}\,d\lambda.$
For $|\lambda|>\eta$, the density $f_{H}$ is bounded, and thus $g_{n}(\lambda)$ is uniformly bounded in $n$. Consequently, $I_{0,n}^{(\infty)}=O(1)$. Since $p>1/2$, we have $\Delta_{n}^{1-2p}\to\infty$, which implies $I_{0,n}^{(\infty)}=o(\Delta_{n}^{1-2p})$.
For the low-frequency part $I_{0,n}^{(0)}$, we use the change of variables $\lambda=\Delta_{n}x$:
$I_{0,n}^{(0)}=\Delta_{n}\int_{|x|\leq\eta/\Delta_{n}}\left(\frac{f_{H}(\Delta_{n}x)}{1+\gamma_{n}f_{H}(\Delta_{n}x)}\right)^{2}dx.$
By (31), uniformly on compact sets of $x\neq 0$, we have $f_{H}(\Delta_{n}x)=c_{H}\Delta_{n}^{-p}|x|^{-p}(1+o(1))$. Moreover, $\gamma_{n}f_{H}(\Delta_{n}x)=\sigma^{2}\Delta_{n}^{p}\cdot c_{H}\Delta_{n}^{-p}|x|^{-p}(1+o(1))=\sigma^{2}c_{H}|x|^{-p}(1+o(1))$. Thus, for each fixed $x\neq 0$,
$\left(\frac{f_{H}(\Delta_{n}x)}{1+\gamma_{n}f_{H}(\Delta_{n}x)}\right)^{2}=\Delta_{n}^{-2p}w(x)\,(1+o(1)).$
Applying Lemma 4 (with $b=c=f_{H}$) and the change of variables $\lambda=\Delta_{n}x$, the integrand is dominated by an integrable function of the form $C\Delta_{n}^{-2p}w(x)$. Using the Dominated Convergence Theorem and the fact that $\eta/\Delta_{n}\to\infty$, we obtain
$I_{0,n}^{(0)}=\Delta_{n}^{1-2p}\int_{\mathbb{R}}w(x)\,dx+o\bigl(\Delta_{n}^{1-2p}\bigr)=\Delta_{n}^{1-2p}\bigl(J_{0}+o(1)\bigr).$
Combining this with the estimate for $I_{0,n}^{(\infty)}$ yields $I_{0,n}=\Delta_{n}^{1-2p}(J_{0}+o(1))$. Inserting this into (21) completes the proof. ∎
Now we can deal with the part of the parameter $\sigma$ of the score funtion
Proposition 3.
For $S_{\sigma,n}=\partial_{\sigma}\ell_{n}$ and $\mathcal{T}_{n}=n\Delta_{n}\rightarrow\infty$, we have the central limit theorem as $n\rightarrow\infty$:
$\frac{S_{\sigma,n}}{\sqrt{\mathcal{T}_{n}}}\ \Rightarrow\ \mathcal{N}\!\left(0,\ \frac{\sigma^{2}}{\pi}J_{0}(H,\sigma)\right),$ (34)
with $J_{0}(H,\sigma)$ defined in (28).
Proof.
First of all we will Verify the condition $\left\lVert C_{n}\right\rVert_{\mathrm{op}}/\left\lVert C_{n}\right\rVert_{F}\to 0$. As we know $T_{n}$ is positive semidefinite (as a Toeplitz covariance matrix), hence $A_{n}$ is positive definite. Consequently, the eigenvalues of
$C_{n}=A_{n}^{-1/2}T_{n}A_{n}^{-1/2}$
are of the form $\lambda/(1+\gamma_{n}\lambda)$ where $\lambda\geq 0$ are the eigenvalues of $T_{n}$. The function $x\mapsto x/(1+\gamma_{n}x)$ is increasing on $0,\infty)$ and bounded by $1/\gamma_{n}$. Thus,
$\left\lVert C_{n}\right\rVert_{\mathrm{op}}=\max_{\lambda\in\text{spec}(T_{n})}\frac{\lambda}{1+\gamma_{n}\lambda}\leq\frac{1}{\gamma_{n}}.$
Using the trace asymptotic from Lemma 6, we have
$\left\lVert C_{n}\right\rVert_{F}=\sqrt{\mathrm{tr}(C_{n}^{2})}=\sqrt{\frac{n}{2\pi}\Delta_{n}^{1-2p}\bigl(J_{0}+o(1)\bigr)}.$
Recalling that $\gamma_{n}=\sigma^{2}\Delta_{n}^{p}$, the ratio behaves as:
$\frac{\left\lVert C_{n}\right\rVert_{\mathrm{op}}}{\left\lVert C_{n}\right\rVert_{F}}\leq\frac{1/\gamma_{n}}{\left\lVert C_{n}\right\rVert_{F}}\asymp\frac{\Delta_{n}^{-p}}{\sqrt{n}\Delta_{n}^{\frac{1-2p}{2}}}=\frac{1}{\sqrt{n\Delta_{n}}}=\frac{1}{\sqrt{\mathcal{T}_{n}}}\to 0,$
since $\mathcal{T}_{n}=n\Delta_{n}=n^{1-\alpha}\to\infty$ for $\alpha\in(0,1)$.
Now, set $Q_{n}(C_{n}):=Z^{\top}C_{n}Z-\mathrm{tr}(C_{n})$. By Lemma 2,
$\frac{Q_{n}(C_{n})}{\sqrt{2}\left\lVert C_{n}\right\rVert_{F}}\Rightarrow\mathcal{N}(0,1).$
From Lemma 1, $S_{\sigma,n}=\frac{\gamma_{n}}{\sigma}Q_{n}(C_{n})$. Thus,
$\frac{S_{\sigma,n}}{\sqrt{2}\,\frac{\gamma_{n}}{\sigma}\left\lVert C_{n}\right\rVert_{F}}=\frac{Q_{n}(C_{n})}{\sqrt{2}\left\lVert C_{n}\right\rVert_{F}}\Rightarrow\mathcal{N}(0,1).$
It remains to identify the variance scale. Using $\gamma_{n}=\sigma^{2}\Delta_{n}^{p}$ and the asymptotic for $\left\lVert C_{n}\right\rVert_{F}^{2}=\mathrm{tr}(C_{n}^{2})$, we obtain
$\left(\sqrt{2}\,\frac{\gamma_{n}}{\sigma}\left\lVert C_{n}\right\rVert_{F}\right)^{2}=2\frac{\sigma^{4}\Delta_{n}^{2p}}{\sigma^{2}}\cdot\frac{n}{2\pi}\Delta_{n}^{1-2p}\Big(J_{0}+o(1)\Big)=\frac{\sigma^{2}}{\pi}\,(n\Delta_{n})\Big(J_{0}+o(1)\Big)=\frac{\sigma^{2}}{\pi}\,\mathcal{T}_{n}\Big(J_{0}+o(1)\Big).$
Therefore,
$\frac{S_{\sigma,n}}{\sqrt{\mathcal{T}_{n}}}\Rightarrow\mathcal{N}\!\left(0,\ \frac{\sigma^{2}}{\pi}J_{0}\right),$
which yields (34). ∎
3.3 The CLT for the Remainder of Score function of $H$
In Propostion 1 we know that $S_{H,n}$ has two part, one is associated with $S_{\sigma,H}$ and the other one connected with matrix $D_{n}$ we called the remainder one. To get the CLT of this remainder part we will use the same way as in $S_{\sigma,H}$.
Lemma 7.
Recall the matrices defined in Lemma 1:
$A_{n}:=I_{n}+\gamma_{n}T_{n}(f_{H}),\qquad C_{n}:=A_{n}^{-1/2}T_{n}(f_{H})A_{n}^{-1/2},\qquad D_{n}:=A_{n}^{-1/2}T_{n}(\dot{f}_{H})A_{n}^{-1/2}.$
Then, as $n\to\infty$,
$\left\lVert D_{n}\right\rVert_{\mathrm{op}}=O\!\big(\gamma_{n}^{-1}\log n\big).$ (35)
Proof.
Still with $T_{n}:=T_{n}(f_{H})$ and $\dot{T}_{n}:=T_{n}(\dot{f}_{H})$ we define the auxiliary matrices
$B_{n}:=T_{n}^{1/2}A_{n}^{-1/2}=A_{n}^{-1/2}T_{n}^{1/2},\qquad E_{n}:=T_{n}^{-1/2}\dot{T}_{n}\,T_{n}^{-1/2}.$
Then we can factorize $D_{n}$ as $D_{n}=B_{n}E_{n}B_{n}$. Consequently,
$\left\lVert D_{n}\right\rVert_{\mathrm{op}}\leq\left\lVert B_{n}\right\rVert_{\mathrm{op}}^{2}\,\left\lVert E_{n}\right\rVert_{\mathrm{op}}.$
First, observe that
$\left\lVert B_{n}\right\rVert_{\mathrm{op}}^{2}=\left\lVert B_{n}^{2}\right\rVert_{\mathrm{op}}=\left\lVert A_{n}^{-1/2}T_{n}A_{n}^{-1/2}\right\rVert_{\mathrm{op}}=\left\lVert C_{n}\right\rVert_{\mathrm{op}}.$
As shown in the proof of Proposition 3, the eigenvalues of $C_{n}$ are bounded by $1/\gamma_{n}$. Thus,
$\left\lVert D_{n}\right\rVert_{\mathrm{op}}\leq\frac{1}{\gamma_{n}}\,\left\lVert E_{n}\right\rVert_{\mathrm{op}}.$ (36)
It remains to show $\left\lVert E_{n}\right\rVert_{\mathrm{op}}=O(\log n)$. Define the log-derivative symbol
$b_{H}(\lambda):=\partial_{H}\log f_{H}(\lambda)=\frac{\dot{f}_{H}(\lambda)}{f_{H}(\lambda)}.$
By Lemma 3, we have $f_{H}(\lambda)=|2\sin(\lambda/2)|^{-p}\ell_{H}(\lambda)$ with $\ell_{H}$ strictly positive and $C^{1}$. Therefore,
$b_{H}(\lambda)=\partial_{H}\log\ell_{H}(\lambda)\;-\;2\log\bigl(2|\sin(\lambda/2)|\bigr).$
The first term is bounded with absolutely summable Fourier coefficients. The second term involves the Fourier series of the log-sine function:
$\log\bigl(2|\sin(\lambda/2)|\bigr)=\sum_{k=1}^{\infty}\frac{\cos(k\lambda)}{k},\qquad\lambda\in(-\pi,\pi).$
The Fourier coefficients $(b_{H})_{k}$ satisfy $|(b_{H})_{k}|\leq C/|k|$ for $k\neq 0$. Hence, the operator norm of the Toeplitz matrix generated by $b_{H}$ satisfies
$\left\lVert T_{n}(b_{H})\right\rVert_{\mathrm{op}}\leq\sum_{k=-(n-1)}^{n-1}|(b_{H})_{k}|\leq C^{\prime}\sum_{k=1}^{n}\frac{1}{k}\leq C^{\prime\prime}\log n.$
Finally, $E_{n}$ approximates $T_{n}(b_{H})$. Specifically, $E_{n}=T_{n}^{-1/2}T_{n}(b_{H}f_{H})T_{n}^{-1/2}$.
By standard preconditioning results for Toeplitz matrices, the sandwiched matrix $E_{n}=T_{n}^{-1/2}T_{n}(\dot{f}_{H})T_{n}^{-1/2}$ admits the decomposition $E_{n}=T_{n}(b_{H})+R_{n}$ with $\left\lVert R_{n}\right\rVert_{\mathrm{op}}=O(1)$ (see [2] or [5]); Substituting this into (36) yields (35). ∎
Lemma 8.
Under (31), (32) and (23), we have
$\mathrm{tr}(D_{n}^{2})=\frac{n}{2\pi}\Delta_{n}^{1-2p}\Bigl(4\ln^{2}(1/\Delta_{n})\,J_{0}(H,\sigma)+4\ln(1/\Delta_{n})\,J_{1}(H,\sigma)+J_{2}(H,\sigma)+o\!\big(\ln^{2}(1/\Delta_{n})\big)\Bigr).$ (37)
Proof.
By (23), it suffices to analyze
$I_{2,n}:=\int_{-\pi}^{\pi}h_{n}(\lambda)^{2}\,d\lambda=\int_{-\pi}^{\pi}\frac{\dot{f}_{H}(\lambda)^{2}}{(1+\gamma_{n}f_{H}(\lambda))^{2}}\,d\lambda.$
Fix $\eta\in(0,\pi)$ and split $I_{2,n}=I_{2,n}^{(0)}+I_{2,n}^{(\infty)}$ with
$I_{2,n}^{(0)}:=\int_{|\lambda|\leq\eta}h_{n}(\lambda)^{2}\,d\lambda,\qquad I_{2,n}^{(\infty)}:=\int_{\eta<|\lambda|\leq\pi}h_{n}(\lambda)^{2}\,d\lambda.$
On $\{\eta<|\lambda|\leq\pi\}$ both $f_{H}$ and $\dot{f}_{H}$ are bounded and $a_{n}(\lambda)\geq 1$, hence $h_{n}(\lambda)^{2}\leq C_{\eta}$ uniformly in $n$, so $I_{2,n}^{(\infty)}=O(1)$. Since $\Delta_{n}^{1-2p}\ln^{2}(1/\Delta_{n})\to\infty$ for $p>1/2$,
$I_{2,n}^{(\infty)}=o\!\big(\Delta_{n}^{1-2p}\ln^{2}(1/\Delta_{n})\big).$
With $\lambda=\Delta_{n}x$,
$I_{2,n}^{(0)}=\Delta_{n}\int_{|x|\leq\eta/\Delta_{n}}\frac{\dot{f}_{H}(\Delta_{n}x)^{2}}{(1+\gamma_{n}f_{H}(\Delta_{n}x))^{2}}\,dx.$
Write again
$\kappa_{n}(x):=\frac{f_{H}(\Delta_{n}x)^{2}}{(1+\gamma_{n}f_{H}(\Delta_{n}x))^{2}},\qquad b_{n}(x):=\frac{\dot{f}_{H}(\Delta_{n}x)}{f_{H}(\Delta_{n}x)}.$
Then
$I_{2,n}^{(0)}=\Delta_{n}\int_{|x|\leq\eta/\Delta_{n}}\kappa_{n}(x)\,b_{n}(x)^{2}\,dx.$ (38)
Now, for each fixed $x\neq 0$,
$\kappa_{n}(x)=\Delta_{n}^{-2p}w(x)\bigl(1+o(1)\bigr),$
and
$b_{n}(x)=2\ln(1/\Delta_{n})+\bigl(C_{H}-2\ln|x|\bigr)+o(1).$
Therefore,
$b_{n}(x)^{2}$ $=\Bigl(2\ln(1/\Delta_{n})+C_{H}-2\ln|x|+o(1)\Bigr)^{2}$
$=4\ln^{2}(1/\Delta_{n})+4\ln(1/\Delta_{n})\bigl(C_{H}-2\ln|x|\bigr)+\bigl(C_{H}-2\ln|x|\bigr)^{2}+o\bigl(\ln^{2}(1/\Delta_{n})\bigr).$ (39)
Multiplying by $\Delta_{n}\,\kappa_{n}(x)$ gives the pointwise expansion
$\Delta_{n}\,\kappa_{n}(x)b_{n}(x)^{2}$ $=\Delta_{n}^{1-2p}w(x)\Bigl(4\ln^{2}(1/\Delta_{n})+4\ln(1/\Delta_{n})\bigl(C_{H}-2\ln|x|\bigr)+\bigl(C_{H}-2\ln|x|\bigr)^{2}\Bigr)$
$\quad+o\!\big(\Delta_{n}^{1-2p}\ln^{2}(1/\Delta_{n})\big).$ (40)
In Lemma 4 with $b=c=\dot{f}_{H}$ yields, for $|\lambda|\leq\eta$,
$\frac{\dot{f}_{H}(\lambda)^{2}}{(1+\gamma_{n}f_{H}(\lambda))^{2}}\leq C\,\frac{f_{H}(\lambda)^{2}\bigl(1+|\log|\lambda||\bigr)^{2}}{(1+\gamma_{n}f_{H}(\lambda))^{2}}.$
With $\lambda=\Delta_{n}x$ and the same use of (31) as before, we obtain for $n$ large
$\Delta_{n}\,\kappa_{n}(x)b_{n}(x)^{2}\leq C^{\prime}\,\Delta_{n}^{1-2p}\,w(x)\Bigl(1+|\log\Delta_{n}|+|\log|x||\Bigr)^{2}.$
By your assumption right after (27), we have $w\in L^{1}(\mathbb{R})$ and $w(x)\,|\log|x||^{2}\in L^{1}(\mathbb{R})$. Thus the right-hand side is integrable in $x$ (again up to the factor $\Delta_{n}^{1-2p}$), so dominated convergence applies to (38), and we may extend the truncation $|x|\leq\eta/\Delta_{n}$ to $\mathbb{R}$ because $\eta/\Delta_{n}\to\infty$.
Consequently, integrating the expansion (40) yields
$I_{2,n}^{(0)}=\Delta_{n}^{1-2p}\Bigl($ $4\ln^{2}(1/\Delta_{n})\int_{\mathbb{R}}w(x)\,dx+4\ln(1/\Delta_{n})\int_{\mathbb{R}}w(x)\bigl(C_{H}-2\ln|x|\bigr)\,dx$
$+\int_{\mathbb{R}}w(x)\bigl(C_{H}-2\ln|x|\bigr)^{2}\,dx+o\bigl(\ln^{2}(1/\Delta_{n})\bigr)\Bigr).$
By (28)–(30), this becomes
$I_{2,n}^{(0)}=\Delta_{n}^{1-2p}\Bigl(4\ln^{2}(1/\Delta_{n})\,J_{0}(H,\sigma)+4\ln(1/\Delta_{n})\,J_{1}(H,\sigma)+J_{2}(H,\sigma)+o\bigl(\ln^{2}(1/\Delta_{n})\bigr)\Bigr).$
Combining with $I_{2,n}^{(\infty)}=O(1)$ gives the same expansion for $I_{2,n}$. Finally insert into (23) to obtain (37). ∎
Proposition 4.
Let
$R_{H,n}:=\frac{\gamma_{n}}{2}\Big(Z^{\top}D_{n}Z-\mathrm{tr}(D_{n})\Big),\qquad Z\sim\mathcal{N}(0,I_{n}),$
with $D_{n}=A_{n}^{-1/2}T_{n}(\dot{f}_{H})A_{n}^{-1/2}$, $\Delta_{n}=n^{-\alpha}$, $\alpha\in(0,1)$, $\gamma_{n}=\sigma^{2}\Delta_{n}^{p}$, $p=2H-1\in(1/2,1)$, and $\mathcal{T}_{n}:=n\Delta_{n}\to\infty$. With (37) and Lemma 7 we have
$\frac{R_{H,n}}{|\ln\Delta_{n}|\,\sqrt{\mathcal{T}_{n}}}\ \Rightarrow\ \mathcal{N}\!\left(0,\ \frac{\sigma^{4}}{\pi}J_{0}(H,\sigma)\right).$ (41)
Proof.
Set $Q_{n}(D_{n}):=Z^{\top}D_{n}Z-\mathrm{tr}(D_{n})$. By (37), we have
$\left\lVert D_{n}\right\rVert_{F}^{2}=\mathrm{tr}(D_{n}^{2})\asymp n\Delta_{n}^{1-2p}\ln^{2}(1/\Delta_{n}).$
By Lemma 7, $\left\lVert D_{n}\right\rVert_{\mathrm{op}}=O(\gamma_{n}^{-1}\log n)$, hence
$\frac{\left\lVert D_{n}\right\rVert_{\mathrm{op}}}{\left\lVert D_{n}\right\rVert_{F}}=O\!\left(\frac{1}{\sqrt{n\Delta_{n}}}\right)=O\!\left(\frac{1}{\sqrt{\mathcal{T}_{n}}}\right)\to 0.$
Therefore Lemma 2 yields
$\frac{Q_{n}(D_{n})}{\sqrt{2}\left\lVert D_{n}\right\rVert_{F}}\Rightarrow\mathcal{N}(0,1).$
Since $R_{H,n}=\frac{\gamma_{n}}{2}Q_{n}(D_{n})$, we get
$\frac{R_{H,n}}{\gamma_{n}\left\lVert D_{n}\right\rVert_{F}}\Rightarrow\mathcal{N}\!\left(0,\frac{1}{2}\right).$
Finally, using $\gamma_{n}^{2}=\sigma^{4}\Delta_{n}^{2p}$ and the leading term in (37),
$\gamma_{n}^{2}\left\lVert D_{n}\right\rVert_{F}^{2}=\sigma^{4}\Delta_{n}^{2p}\cdot\frac{n}{2\pi}\Delta_{n}^{1-2p}\Bigl(4\ln^{2}(1/\Delta_{n})\,J_{0}+o(\ln^{2}(1/\Delta_{n}))\Bigr)=\frac{2\sigma^{4}}{\pi}\,\mathcal{T}_{n}\,\ln^{2}(1/\Delta_{n})\Bigl(J_{0}+o(1)\Bigr),$
which implies (41). ∎
Remark 2.
The finer terms involving $J_{1}$ and $J_{2}$ in (37) will be used later (Section 4) for the orthogonalization step and for identifying the full Fisher information matrix. The explicit scaling in (41) only needs the leading $\ln^{2}(1/\Delta_{n})$ contribution, hence only $J_{0}$ appears in the variance.
4 CLT for the projected score vector
In Appendix B we know that $S_{\sigma,n}$ and the remainder of $H$ (Standard quadratic formula of $C_{n}$ and $D_{n}$) are strongly dependent, the correlation coefficient tends to $1$ when $n\rightarrow\infty$. This result is far from the LAN property with non-singular matrix. In order to avoid this problem, we take the idea for the projection from $D_{n}$ to $C_{n}$.
4.1 Projection and exact orthogonality
First, we give the exact definition of the projection:
Definition 2.
Let us define the deterministic projection coefficient
$\mathbf{a}_{n}:=\frac{\mathrm{tr}(C_{n}D_{n})}{\mathrm{tr}(C_{n}^{2})},$ (42)
also the orthogonalized matrix $D_{n}^{\perp}:=D_{n}-\mathbf{a}_{n}C_{n}$ and the orthogonalized remainder
$R_{H,n}^{\perp}:=\frac{\gamma_{n}}{2}\Big(Z^{\top}D_{n}^{\perp}Z-\mathrm{tr}(D_{n}^{\perp})\Big)=R_{H,n}-\frac{\sigma\mathbf{a}_{n}}{2}\,S_{\sigma,n}.$ (43)
By construction, $\mathrm{tr}(C_{n}D_{n}^{\perp})=0$.
Lemma 9.
With $\mathbf{a}_{n}$ and $R_{H,n}^{\perp}$ defined above,
$\mathrm{Cov}(S_{\sigma,n},R_{H,n}^{\perp})=0.$
Proof.
Write $S_{\sigma,n}=\frac{\gamma_{n}}{\sigma}(Z^{\top}C_{n}Z-\mathrm{tr}C_{n})$ and $R_{H,n}^{\perp}=\frac{\gamma_{n}}{2}(Z^{\top}D_{n}^{\perp}Z-\mathrm{tr}D_{n}^{\perp})$. By Wick’s identity for centered Gaussian quadratic forms,
$\mathrm{Cov}(S_{\sigma,n},R_{H,n}^{\perp})=\frac{\gamma_{n}}{\sigma}\cdot\frac{\gamma_{n}}{2}\cdot 2\mathrm{tr}(C_{n}D_{n}^{\perp})=\frac{\gamma_{n}^{2}}{\sigma}\Big(\mathrm{tr}(C_{n}D_{n})-\mathbf{a}_{n}\mathrm{tr}(C_{n}^{2})\Big)=0,$
by the definition of $\mathbf{a}_{n}$. ∎
4.2 Asymptotics of the projection coefficient and the orthogonal Frobenius norm
Before the expansion of $\mathbf{a}_{n}$, we first construct the trace approximation of the product $C_{n}D_{n}$:
Lemma 10.
Under the expansions (31), (32) and the approximation (22), we have
$\mathrm{tr}(C_{n}D_{n})=\frac{n}{2\pi}\Delta_{n}^{1-2p}\Bigl(2\ln(1/\Delta_{n})\,J_{0}(H,\sigma)+J_{1}(H,\sigma)+o\!\big(\ln(1/\Delta_{n})\big)\Bigr),$ (44)
where $J_{0}$ and $J_{1}$ are defined in (28) and (29).
Proof.
The proof is almost the same as in Lemma 8. By (22), it suffices to analyze the integral
$I_{1,n}:=\int_{-\pi}^{\pi}g_{n}(\lambda)h_{n}(\lambda)\,d\lambda=\int_{-\pi}^{\pi}\frac{f_{H}(\lambda)\dot{f}_{H}(\lambda)}{(1+\gamma_{n}f_{H}(\lambda))^{2}}\,d\lambda.$
Fix a small $\eta\in(0,\pi)$ and split $I_{1,n}=I_{1,n}^{(0)}+I_{1,n}^{(\infty)}$ as in the proof of Lemma 6.
On the domain $\{\eta<|\lambda|\leq\pi\}$, the functions $f_{H}$ and $\dot{f}_{H}$ are bounded, and $a_{n}(\lambda)\geq 1$. Thus, $|g_{n}(\lambda)h_{n}(\lambda)|\leq C_{\eta}$ uniformly in $n$, implying $I_{1,n}^{(\infty)}=O(1)$. Since $p>1/2$, $\Delta_{n}^{1-2p}\to\infty$, so $I_{1,n}^{(\infty)}=o(\Delta_{n}^{1-2p}\ln(1/\Delta_{n}))$.
On $|\lambda|\leq\eta$, use the change of variables $\lambda=\Delta_{n}x$:
$I_{1,n}^{(0)}=\Delta_{n}\int_{|x|\leq\eta/\Delta_{n}}\frac{f_{H}(\Delta_{n}x)\,\dot{f}_{H}(\Delta_{n}x)}{\bigl(1+\gamma_{n}f_{H}(\Delta_{n}x)\bigr)^{2}}\,dx.$
Define the auxiliary functions
$\kappa_{n}(x):=\frac{f_{H}(\Delta_{n}x)^{2}}{\bigl(1+\gamma_{n}f_{H}(\Delta_{n}x)\bigr)^{2}},\qquad b_{n}(x):=\frac{\dot{f}_{H}(\Delta_{n}x)}{f_{H}(\Delta_{n}x)}.$
Then the integral becomes
$I_{1,n}^{(0)}=\Delta_{n}\int_{|x|\leq\eta/\Delta_{n}}\kappa_{n}(x)\,b_{n}(x)\,dx.$ (45)
By (31), for each fixed $x\neq 0$,
$\gamma_{n}f_{H}(\Delta_{n}x)\sim\sigma^{2}c_{H}|x|^{-p},$
which implies (as shown in Lemma 6)
$\kappa_{n}(x)=\Delta_{n}^{-2p}w(x)\bigl(1+o(1)\bigr).$ (46)
For the term $b_{n}(x)$, using (32), we have
$b_{n}(x)=\frac{\dot{f}_{H}(\Delta_{n}x)}{f_{H}(\Delta_{n}x)}=C_{H}-2\ln|\Delta_{n}x|+o(1)=2\ln(1/\Delta_{n})+\bigl(C_{H}-2\ln|x|\bigr)+o(1).$
Combining (46) and the expansion of $b_{n}(x)$, the integrand behaves as
$\Delta_{n}\,\kappa_{n}(x)b_{n}(x)=\Delta_{n}^{1-2p}w(x)\Bigl(2\ln(1/\Delta_{n})+C_{H}-2\ln|x|\Bigr)+o\!\big(\Delta_{n}^{1-2p}\ln(1/\Delta_{n})\big).$ (47)
By Lemma 4, the integrand is dominated by an integrable function involving logarithmic factors. Thus, we can apply the Dominated Convergence Theorem. Integrating the leading terms over $\mathbb{R}$ (since $\eta/\Delta_{n}\to\infty$) yields:
$I_{1,n}^{(0)}$ $=\Delta_{n}^{1-2p}\left[2\ln(1/\Delta_{n})\int_{\mathbb{R}}w(x)dx+\int_{\mathbb{R}}w(x)(C_{H}-2\ln|x|)dx+o(\ln(1/\Delta_{n}))\right]$
$=\Delta_{n}^{1-2p}\Bigl(2\ln(1/\Delta_{n})J_{0}+J_{1}+o(\ln(1/\Delta_{n}))\Bigr).$
Combining this with $I_{1,n}^{(\infty)}$ and inserting into (22) proves (44). ∎
Now, we give the second-order expansion of $\mathbf{a}_{n}$:
Lemma 11.
Assume the trace asymptotics (33) and (44). Then
$\mathbf{a}_{n}=\frac{\mathrm{tr}(C_{n}D_{n})}{\mathrm{tr}(C_{n}^{2})}=2\ln\!\Big(\frac{1}{\Delta_{n}}\Big)+m(H,\sigma)+o(1),\qquad m(H,\sigma)=\frac{J_{1}(H,\sigma)}{J_{0}(H,\sigma)}.$ (48)
Proof.
By (42) and (33)–(44),
$\mathbf{a}_{n}=\frac{2\ln(1/\Delta_{n})\,J_{0}+J_{1}+o(1)}{J_{0}+o(1)}=2\ln\!\Big(\frac{1}{\Delta_{n}}\Big)+\frac{J_{1}}{J_{0}}+o(1).$
We will obtain the Frobenius norm of the orthogonalized matrix:
Lemma 12.
Let $D_{n}^{\perp}=D_{n}-\mathbf{a}_{n}C_{n}$ with $\mathbf{a}_{n}$ defined by (42). Then
$\|D_{n}^{\perp}\|_{F}^{2}=\mathrm{tr}\big((D_{n}^{\perp})^{2}\big)=\frac{n}{2\pi}\Delta_{n}^{1-2p}\Big(J_{\perp}+o(1)\Big),\qquad J_{\perp}:=J_{2}-\frac{J_{1}^{2}}{J_{0}}\in(0,\infty).$ (49)
Proof.
Expand
$\mathrm{tr}\big((D_{n}-\mathbf{a}_{n}C_{n})^{2}\big)=\mathrm{tr}(D_{n}^{2})-2\mathbf{a}_{n}\mathrm{tr}(C_{n}D_{n})+\mathbf{a}_{n}^{2}\mathrm{tr}(C_{n}^{2}).$
Insert (33), (44), (37), and use $\mathbf{a}_{n}=2\ln(1/\Delta_{n})+m+o(1)$ with $m=J_{1}/J_{0}$ from Lemma 11. A direct algebraic cancellation shows that the coefficients of $\ln^{2}(1/\Delta_{n})$ and $\ln(1/\Delta_{n})$ vanish, and the remaining constant term equals $J_{2}-m^{2}J_{0}=J_{2}-J_{1}^{2}/J_{0}$. This yields (49). Positivity follows since $J_{\perp}=\int_{\mathbb{R}}w(x)(q(x)-m)^{2}dx$ with $q(x)=C_{H}-2\ln|x|$ not $w$-a.e. constant. ∎
Now, we will try to get the structure of $\|\cdot\|_{\mathrm{op}}/\|\cdot\|_{F}\to 0$ for $D_{n}^{\perp}$.
Lemma 13.
Fix $H\in(3/4,1)$ and set $p:=2H-1\in(1/2,1)$. Let $\Delta_{n}=n^{-\alpha}$ with $\alpha\in(0,1)$, $\gamma_{n}=\sigma^{2}\Delta_{n}^{p}$ and $\mathcal{T}_{n}=n\Delta_{n}\to\infty$. Recall $D_{n}^{\perp}=D_{n}-\mathbf{a}_{n}C_{n}$ with $\mathbf{a}_{n}$ defined by (42). Then
$\frac{\left\lVert D_{n}^{\perp}\right\rVert_{\mathrm{op}}}{\left\lVert D_{n}^{\perp}\right\rVert_{F}}\longrightarrow 0.$
Proof.
From Lemma 7 we have $\left\lVert C_{n}\right\rVert_{\mathrm{op}}\leq 1/\gamma_{n}$ and $\left\lVert D_{n}\right\rVert_{\mathrm{op}}=O\big(\gamma_{n}^{-1}\log n\big)$. Now, by the triangle inequality,
$\left\lVert D_{n}^{\perp}\right\rVert_{\mathrm{op}}\leq\left\lVert D_{n}\right\rVert_{\mathrm{op}}+|\mathbf{a}_{n}|\,\left\lVert C_{n}\right\rVert_{\mathrm{op}}\leq\left\lVert D_{n}\right\rVert_{\mathrm{op}}+\frac{|\mathbf{a}_{n}|}{\gamma_{n}}.$
From Lemma 11 we have $\mathbf{a}_{n}=2\ln(1/\Delta_{n})+O(1)$, hence $|\mathbf{a}_{n}|=O(\log n)$ because $\Delta_{n}=n^{-\alpha}$. Together with Lemma 7 this yields
$\left\lVert D_{n}^{\perp}\right\rVert_{\mathrm{op}}=O\!\Big(\frac{\log n}{\gamma_{n}}\Big).$
On the other hand, Lemma 12 gives
$\left\lVert D_{n}^{\perp}\right\rVert_{F}^{2}=\frac{n}{2\pi}\Delta_{n}^{1-2p}\big(J_{\perp}+o(1)\big),\qquad J_{\perp}>0,$
so $\left\lVert D_{n}^{\perp}\right\rVert_{F}\asymp\sqrt{n}\Delta_{n}^{\frac{1-2p}{2}}$. Since $\gamma_{n}=\sigma^{2}\Delta_{n}^{p}$, there exists a constant $C>0$ such that
$\frac{\left\lVert D_{n}^{\perp}\right\rVert_{\mathrm{op}}}{\left\lVert D_{n}^{\perp}\right\rVert_{F}}\leq C\frac{\Delta_{n}^{-p}\log n}{\sqrt{n}\,\Delta_{n}^{\frac{1-2p}{2}}}=C\frac{\log n}{\sqrt{n\Delta_{n}}}=C\frac{\log n}{\sqrt{\mathcal{T}_{n}}}\longrightarrow 0,$
because $\mathcal{T}_{n}=n\Delta_{n}\to\infty$. ∎
4.3 Joint CLT after multiplying by the rate matrix
Define the lower-triangular matrices
$M_{n}^{(1)}:=\begin{pmatrix}1&0\\ -\sigma\ln(\Delta_{n})&1\end{pmatrix},\qquad M_{n}^{(2)}:=\begin{pmatrix}1&0\\ -\frac{\sigma\mathbf{a}_{n}}{2}&1\end{pmatrix},\qquad M_{n}:=M_{n}^{(2)}M_{n}^{(1)}=\begin{pmatrix}1&0\\ -\sigma\ln(\Delta_{n})-\frac{\sigma\mathbf{a}_{n}}{2}&1\end{pmatrix}$. (50)
Recall from Lemma 1 that $S_{H,n}=\sigma\ln(\Delta_{n})\,S_{\sigma,n}+R_{H,n}$. Then
$M_{n}\binom{S_{\sigma,n}}{S_{H,n}}=\binom{S_{\sigma,n}}{R_{H,n}^{\perp}}$.
Proposition 5.
Assume $\mathcal{T}_{n}=n\Delta_{n}\to\infty$. Define
$\Xi_{n}:=\frac{1}{\sqrt{\mathcal{T}_{n}}}\,M_{n}\binom{S_{\sigma,n}}{S_{H,n}}=\binom{\frac{S_{\sigma,n}}{\sqrt{\mathcal{T}_{n}}}}{\frac{R_{H,n}^{\perp}}{\sqrt{\mathcal{T}_{n}}}}$.
Then, as $n\to\infty$,
$\Xi_{n}\ \Rightarrow\ \mathcal{N}\!\left(0,\ I^{\perp}\right),\qquad I^{\perp}:=\begin{pmatrix}\frac{\sigma^{2}}{\pi}J_{0}(H,\sigma)&0\\ 0&\frac{\sigma^{4}}{4\pi}J_{\perp}(H,\sigma)\end{pmatrix},\qquad J_{\perp}=J_{2}-\frac{J_{1}^{2}}{J_{0}}>0.$
Proof.
By Proposition 3,
$\frac{S_{\sigma,n}}{\sqrt{\mathcal{T}_{n}}}\Rightarrow\mathcal{N}\!\left(0,\frac{\sigma^{2}}{\pi}J_{0}\right).$
Moreover,
$R_{H,n}^{\perp}=\frac{\gamma_{n}}{2}\Big(Z^{\top}D_{n}^{\perp}Z-\mathrm{tr}(D_{n}^{\perp})\Big).$
By Lemma 13 and Lemma 2,
$\frac{Z^{\top}D_{n}^{\perp}Z-\mathrm{tr}(D_{n}^{\perp})}{\sqrt{2}\left\lVert D_{n}^{\perp}\right\rVert_{F}}\Rightarrow\mathcal{N}(0,1).$
Using $\left\lVert D_{n}^{\perp}\right\rVert_{F}^{2}=\frac{n}{2\pi}\Delta_{n}^{1-2p}(J_{\perp}+o(1))$ and $\gamma_{n}=\sigma^{2}\Delta_{n}^{p}$ yields
$\frac{R_{H,n}^{\perp}}{\sqrt{\mathcal{T}_{n}}}\Rightarrow\mathcal{N}\!\left(0,\frac{\sigma^{4}}{4\pi}J_{\perp}\right).$
Now, we consider the covariance part. For any fixed $u=(u_{1},u_{2})\in\mathbb{R}^{2}$,
$u_{1}\frac{S_{\sigma,n}}{\sqrt{\mathcal{T}_{n}}}+u_{2}\frac{R_{H,n}^{\perp}}{\sqrt{\mathcal{T}_{n}}}=\frac{\gamma_{n}}{\sqrt{\mathcal{T}_{n}}}\Big(Z^{\top}M_{n,u}Z-\mathrm{tr}(M_{n,u})\Big),\qquad M_{n,u}:=u_{1}\sigma^{-1}C_{n}+u_{2}\tfrac{1}{2}D_{n}^{\perp}.$
Since $\mathrm{tr}(C_{n}D_{n}^{\perp})=0$ by construction, we have
$\left\lVert M_{n,u}\right\rVert_{F}^{2}=u_{1}^{2}\sigma^{-2}\left\lVert C_{n}\right\rVert_{F}^{2}+\frac{u_{2}^{2}}{4}\left\lVert D_{n}^{\perp}\right\rVert_{F}^{2},$
and $\left\lVert M_{n,u}\right\rVert_{\mathrm{op}}\leq|u_{1}|\sigma^{-1}\left\lVert C_{n}\right\rVert_{\mathrm{op}}+|u_{2}|2^{-1}\left\lVert D_{n}^{\perp}\right\rVert_{\mathrm{op}}$. Since $\left\lVert M_{n,u}\right\rVert_{F}^{2}$ is a nonnegative sum of the two terms above, we obtain $\left\lVert M_{n,u}\right\rVert_{\mathrm{op}}/\left\lVert M_{n,u}\right\rVert_{F}\to 0$ from $\left\lVert C_{n}\right\rVert_{\mathrm{op}}/\left\lVert C_{n}\right\rVert_{F}\to 0$ (verified in Proposition 3) and Lemma 13. Applying Lemma 2 to $M_{n,u}$ yields a one-dimensional normal limit for every $u$, hence joint convergence by Cramér–Wold.
Finally, Lemma 9 gives $\mathrm{Cov}(S_{\sigma,n},R_{H,n}^{\perp})=0$ for each $n$, so the limiting covariance is diagonal. ∎
5 LAN Property—Proof of Theorem 2
In this section, we combine the Central Limit Theorem for the projected scores (Proposition 5) with the local expansion of the Gaussian likelihood to prove the main LAN theorem.
5.1 Matrix Taylor Expansions
To establish the Local Asymptotic Normality, we need to analyze the asymptotic behavior of the log-likelihood ratio. This ratio involves non-linear matrix functions of the covariance matrix, specifically the logarithm of the determinant and the matrix inverse. To handle these terms, we rely on second-order Taylor expansions with respect to the perturbation matrix. The following lemma provides the fundamental matrix inequalities required to expand these terms and strictly control the remainder terms using the operator norm and the Frobenius norm (trace of the square).
Lemma 14.
Let $S$ be a real symmetric $n\times n$ matrix with $\left\lVert S\right\rVert_{\mathrm{op}}\leq\frac{1}{2}$, we consider the Log-determinant and inverse of the matrix $I_{n}+S$ and have the following results:
  1. 1.
    $\log\det(I_{n}+S)=\mathrm{tr}(S)-\frac{1}{2}\mathrm{tr}(S^{2})+R_{\log}(S)$, where the remainder satisfies $|R_{\log}(S)|\leq C\,\left\lVert S\right\rVert_{\mathrm{op}}\,\mathrm{tr}(S^{2})$.
  2. 2.
    $(I_{n}+S)^{-1}=I_{n}-S+S^{2}+R_{\mathrm{inv}}(S)\qquad R_{\mathrm{inv}}(S):=\sum_{k\geq 3}(-1)^{k}S^{k}$, and the remainder satisfies
    $\left\lVert R_{\mathrm{inv}}(S)\right\rVert_{\mathrm{op}}\leq C\,\left\lVert S\right\rVert_{\mathrm{op}}^{3},\qquad\mathrm{tr}\!\big(R_{\mathrm{inv}}(S)^{2}\big)\leq C\,\mathrm{tr}(S^{6}).$
Here $C>0$ is a universal constant.
Proof.
Let $\lambda_{1},\dots,\lambda_{n}$ be the eigenvalues of $S$. Since $S$ is symmetric and $\left\lVert S\right\rVert_{\mathrm{op}}\leq 1/2$, we have $|\lambda_{i}|\leq 1/2$ for all $i$.
(1) Log-determinant. For $|x|\leq 1/2$, Taylor’s theorem gives
$\log(1+x)=x-\frac{1}{2}x^{2}+x^{3}\rho(x)\quad\text{with}\quad|\rho(x)|\leq C.$
Summing over eigenvalues yields
$\log\det(I_{n}+S)=\sum_{i=1}^{n}\log(1+\lambda_{i})=\mathrm{tr}(S)-\frac{1}{2}\mathrm{tr}(S^{2})+\sum_{i=1}^{n}\lambda_{i}^{3}\rho(\lambda_{i}).$
Therefore,
$|R_{\log}(S)|\leq C\sum_{i=1}^{n}|\lambda_{i}|^{3}\leq C\,\max_{i}|\lambda_{i}|\sum_{i=1}^{n}\lambda_{i}^{2}\leq C\,\left\lVert S\right\rVert_{\mathrm{op}}\,\mathrm{tr}(S^{2}).$
(2) Inverse. For $|x|\leq 1/2$,
$(1+x)^{-1}=1-x+x^{2}+\sum_{k\geq 3}(-1)^{k}x^{k}.$
By functional calculus this gives
$(I_{n}+S)^{-1}=I_{n}-S+S^{2}+\sum_{k\geq 3}(-1)^{k}S^{k},$
so the stated expression for $R_{\mathrm{inv}}(S)$ holds. Moreover,
$\left\lVert R_{\mathrm{inv}}(S)\right\rVert_{\mathrm{op}}\leq\sum_{k\geq 3}\left\lVert S\right\rVert_{\mathrm{op}}^{k}=\frac{\left\lVert S\right\rVert_{\mathrm{op}}^{3}}{1-\left\lVert S\right\rVert_{\mathrm{op}}}\leq 2\,\left\lVert S\right\rVert_{\mathrm{op}}^{3}.$
Since $R_{\mathrm{inv}}(S)$ is a polynomial in $S$, it commutes with $S$ and we may write $R_{\mathrm{inv}}(S)=S^{3}Q(S)$ with $Q(S):=\sum_{k\geq 0}(-1)^{k}S^{k}=(I_{n}+S)^{-1}$. Hence
$R_{\mathrm{inv}}(S)^{2}=Q(S)^{2}S^{6},\qquad\mathrm{tr}\!\big(R_{\mathrm{inv}}(S)^{2}\big)=\mathrm{tr}\!\big(Q(S)^{2}S^{6}\big)\leq\left\lVert Q(S)\right\rVert_{\mathrm{op}}^{2}\,\mathrm{tr}(S^{6}).$
Finally, $\left\lVert Q(S)\right\rVert_{\mathrm{op}}=\left\lVert(I_{n}+S)^{-1}\right\rVert_{\mathrm{op}}\leq(1-\left\lVert S\right\rVert_{\mathrm{op}})^{-1}\leq 2$, which yields the claimed bound. ∎
5.2 Final Proof of Theorem 2
Fix $h\in\mathbb{R}^{2}$ and set $\theta_{h}:=\theta+hr_{n}^{-1}$. Write $V:=V_{n}(\theta)$ and $V_{h}:=V_{n}(\theta_{h})$. Under $\mathbb{P}_{\theta}$, the whitened vector $Z:=V^{-1/2}X_{n}$ is $\mathcal{N}(0,I_{n})$. Define the symmetric perturbation matrix
$S_{n}(h):=V^{-1/2}(V_{h}-V)V^{-1/2}.$
A standard Gaussian likelihood identity gives the exact log-likelihood ratio representation
$\ell_{n}(\theta_{h})-\ell_{n}(\theta)=-\frac{1}{2}\log\det(I_{n}+S_{n}(h))-\frac{1}{2}\,Z^{\top}\Big((I_{n}+S_{n}(h))^{-1}-I_{n}\Big)Z.$ (51)
First, we verify the condition required for the matrix expansion, namely that the operator norm of the perturbation matrix $S_{n}(h)$ vanishes asymptotically. Recall $V_{n}(\theta)=\Delta_{n}A_{n}(\theta)$ with $A_{n}(\theta):=I_{n}+\gamma_{n}(\theta)T_{n}(H)$, hence $V^{-1/2}=\Delta_{n}^{-1/2}A_{n}(\theta)^{-1/2}$ and
$S_{n}(h)=A_{n}(\theta)^{-1/2}\big(A_{n}(\theta_{h})-A_{n}(\theta)\big)A_{n}(\theta)^{-1/2}.$
Since $r_{n}^{-\top}=(1/\sqrt{\mathcal{T}_{n}})M_{n}$ and, by Lemma 11, the matrix $M_{n}$ is bounded (recall $\mathbf{a}_{n}$ involves $\ln(1/\Delta_{n})$ which cancels with the scaling), we have $\sigma_{h}-\sigma=O(\mathcal{T}_{n}^{-1/2})$ and $H_{h}-H=O(\mathcal{T}_{n}^{-1/2})$.
A first-order Taylor expansion yields
$A_{n}(\theta_{h})-A_{n}(\theta)=(\gamma_{h}-\gamma)T_{n}(H)+\gamma(H_{h}-H)\dot{T}_{n}(H)+\mathcal{R}_{A,n},$
where $\mathcal{R}_{A,n}$ collects the second-order Taylor terms. Moreover, this remainder is negligible after sandwiching: using $\|\ddot{T}_{n}(H)\|_{\mathrm{op}}=O(\log^{2}n)$ (coming from the $\log^{2}|2\sin(\lambda/2)|$ singularity in $\partial_{H}^{2}\log f_{H}$) and $H_{h}-H=O\big(\mathcal{T}_{n}^{-1/2}|\log\Delta_{n}|^{-1}\big)$, we obtain $\|A_{n}(\theta)^{-1/2}\mathcal{R}_{A,n}A_{n}(\theta)^{-1/2}\|_{\mathrm{op}}=o(1)$ (since $|\log\Delta_{n}|\asymp\log n$ and $\mathcal{T}_{n}\to\infty$). Sandwiching by $A_{n}(\theta)^{-1/2}$ gives
$S_{n}(h)=(\gamma_{h}-\gamma)C_{n}(\theta)+\gamma(H_{h}-H)D_{n}(\theta)+R_{S,n},$
with
$C_{n}(\theta)=A_{n}(\theta)^{-1/2}T_{n}(H)A_{n}(\theta)^{-1/2},\,D_{n}(\theta)=A_{n}(\theta)^{-1/2}\dot{T}_{n}(H)A_{n}(\theta)^{-1/2}$
and and $\|R_{S,n}\|_{\mathrm{op}}=o(1)$ uniformly for fixed $h$. Using $\left\lVert C_{n}(\theta)\right\rVert_{\mathrm{op}}\leq 1/\gamma$ and $\left\lVert D_{n}(\theta)\right\rVert_{\mathrm{op}}=O(\gamma^{-1}\log n)$ (by Lemma 7), together with $(\gamma_{h}-\gamma)=O(\gamma\mathcal{T}_{n}^{-1/2})$ and $H_{h}-H=O\big(\mathcal{T}_{n}^{-1/2}|\log\Delta_{n}|^{-1}\big)$, we obtain
$\left\lVert(\gamma_{h}-\gamma)C_{n}(\theta)\right\rVert_{\mathrm{op}}=O(\mathcal{T}_{n}^{-1/2}),\qquad\left\lVert\gamma(H_{h}-H)D_{n}(\theta)\right\rVert_{\mathrm{op}}=O(\mathcal{T}_{n}^{-1/2}\log n),$
and the second-order remainder satisfies $\left\lVert R_{S,n}\right\rVert_{\mathrm{op}}=o(1)$. Since $\mathcal{T}_{n}=n\Delta_{n}=n^{1-\alpha}$ and $\log n/\sqrt{\mathcal{T}_{n}}\to 0$, it follows that $\left\lVert S_{n}(h)\right\rVert_{\mathrm{op}}\to 0$.
Next, having established that $\left\lVert S_{n}(h)\right\rVert_{\mathrm{op}}\to 0$, we apply Lemma 14 to obtain a second-order expansion of the log-likelihood ratio (51). For $n$ large enough we have $\left\lVert S_{n}(h)\right\rVert_{\mathrm{op}}\leq 1/2$, hence
$\ell_{n}(\theta_{h})-\ell_{n}(\theta)$ $=\frac{1}{2}\big(Z^{\top}S_{n}(h)Z-\mathrm{tr}S_{n}(h)\big)-\frac{1}{4}\mathrm{tr}\!\big(S_{n}(h)^{2}\big)$
$\quad-\frac{1}{2}\big(Z^{\top}S_{n}(h)^{2}Z-\mathrm{tr}(S_{n}(h)^{2})\big)-\frac{1}{2}R_{\log}(S_{n}(h))-\frac{1}{2}Z^{\top}R_{\mathrm{inv}}(S_{n}(h))Z.$ (52)
Then, we identify the limits of the principal linear and quadratic terms in this expansion. Let $\delta_{n}:=\theta_{h}-\theta=hr_{n}^{-1}$ (viewed as a vector perturbation). By the Fréchet differentiability of $\theta\mapsto V_{n}(\theta)$,
$S_{n}(h)=\sum_{i=1}^{2}\delta_{n,i}\,M_{i,n}+R_{n},\qquad M_{i,n}:=V^{-1/2}\,\partial_{\theta_{i}}V_{n}(\theta)\,V^{-1/2},$
where the remainder satisfies $\mathrm{tr}(R_{n}^{2})=o(1)$ (the Taylor remainder is quadratic in $\delta_{n}$, and after sandwiching its Frobenius norm is $O(\|\delta_{n}\|^{2})$; since $\|\delta_{n}\|=O(\mathcal{T}_{n}^{-1/2})$, this gives $\mathrm{tr}(R_{n}^{2})=O(\mathcal{T}_{n}^{-1})\to 0$). Therefore,
$\frac{1}{2}\big(Z^{\top}S_{n}(h)Z-\mathrm{tr}S_{n}(h)\big)=\sum_{i=1}^{2}\delta_{n,i}\,\partial_{\theta_{i}}\ell_{n}(\theta)+\frac{1}{2}\big(Z^{\top}R_{n}Z-\mathrm{tr}R_{n}\big).$
The last term is $o_{\mathbb{P}}(1)$ since $\mathrm{Var}(Z^{\top}R_{n}Z-\mathrm{tr}R_{n})=2\mathrm{tr}(R_{n}^{2})=o(1)$. Hence, using $\delta_{n}^{\top}\nabla\ell_{n}(\theta)=h^{\top}(r_{n}^{-\top}\nabla\ell_{n}(\theta))$, we have
$\frac{1}{2}\big(Z^{\top}S_{n}(h)Z-\mathrm{tr}S_{n}(h)\big)=\delta_{n}^{\top}\nabla\ell_{n}(\theta)+o_{\mathbb{P}}(1)=h^{\top}\Xi_{n}+o_{\mathbb{P}}(1),$ (53)
where $\Xi_{n}=r_{n}^{-\,\top}\nabla\ell_{n}(\theta)$.
Next, using $\mathrm{tr}((A+B)^{2})=\mathrm{tr}(A^{2})+2\mathrm{tr}(AB)+\mathrm{tr}(B^{2})$ and Cauchy–Schwarz,
$\mathrm{tr}(S_{n}(h)^{2})=\mathrm{tr}\!\Big(\Big(\sum_{i=1}^{2}\delta_{n,i}M_{i,n}\Big)^{2}\Big)+o(1).$
In the Gaussian covariance model one has the identity (Wick formula)
$\mathrm{Cov}_{\theta}\big(\partial_{\theta_{i}}\ell_{n}(\theta),\partial_{\theta_{j}}\ell_{n}(\theta)\big)=\frac{1}{2}\mathrm{tr}(M_{i,n}M_{j,n})=:\mathcal{I}_{ij,n}(\theta),$
so $\mathcal{I}_{n}(\theta)=(\mathcal{I}_{ij,n}(\theta))_{i,j}$ is both the Fisher information and the covariance of the score. Therefore,
$\frac{1}{4}\mathrm{tr}(S_{n}(h)^{2})=\frac{1}{2}\,\delta_{n}^{\top}\mathcal{I}_{n}(\theta)\,\delta_{n}+o(1)=\frac{1}{2}\,h^{\top}\Big(r_{n}^{-\,\top}\mathcal{I}_{n}(\theta)r_{n}^{-1}\Big)h+o(1).$
Moreover, in the Gaussian covariance model $\mathrm{Cov}_{\theta}(\nabla\ell_{n}(\theta))=\mathcal{I}_{n}(\theta)$, and the same trace asymptotics used in Proposition 5 yield $r_{n}^{-\,\top}\mathcal{I}_{n}(\theta)r_{n}^{-1}\to I^{\perp}$. Consequently,
$\frac{1}{4}\mathrm{tr}(S_{n}(h)^{2})=\frac{1}{2}\,h^{\top}I^{\perp}h+o(1),\qquad\text{in particular }\mathrm{tr}(S_{n}(h)^{2})=O(1).$ (54)
Finally, we control the remaining high-order error terms in (52) and show they converge to zero in probability. Using $\mathrm{Var}(Z^{\top}AZ-\mathrm{tr}A)=2\mathrm{tr}(A^{2})$ and $\mathrm{tr}(S^{4})\leq\left\lVert S\right\rVert_{\mathrm{op}}^{2}\mathrm{tr}(S^{2})$, we obtain
$\mathrm{Var}\big(Z^{\top}S_{n}(h)^{2}Z-\mathrm{tr}(S_{n}(h)^{2})\big)=2\mathrm{tr}(S_{n}(h)^{4})\leq 2\,\left\lVert S_{n}(h)\right\rVert_{\mathrm{op}}^{2}\,\mathrm{tr}(S_{n}(h)^{2})\to 0,$
by the results established above, hence $Z^{\top}S_{n}(h)^{2}Z-\mathrm{tr}(S_{n}(h)^{2})=o_{\mathbb{P}}(1)$. Moreover, Lemma 14 and (54) give
$|R_{\log}(S_{n}(h))|\leq C\,\left\lVert S_{n}(h)\right\rVert_{\mathrm{op}}\,\mathrm{tr}(S_{n}(h)^{2})=o(1).$
Also, write $Z^{\top}R_{\mathrm{inv}}(S_{n}(h))Z=\mathrm{tr}(R_{\mathrm{inv}}(S_{n}(h)))+\big(Z^{\top}R_{\mathrm{inv}}(S_{n}(h))Z-\mathrm{tr}(R_{\mathrm{inv}}(S_{n}(h)))\big)$. Since $R_{\mathrm{inv}}(S)=\sum_{k\geq 3}(-1)^{k}S^{k}$, we have $|\mathrm{tr}(R_{\mathrm{inv}}(S))|\leq\sum_{k\geq 3}|\mathrm{tr}(S^{k})|\leq C\,\left\lVert S\right\rVert_{\mathrm{op}}\mathrm{tr}(S^{2})$, hence $|\mathrm{tr}(R_{\mathrm{inv}}(S_{n}(h)))|=o(1)$. Furthermore,
$\mathrm{Var}\big(Z^{\top}R_{\mathrm{inv}}(S_{n}(h))Z-\mathrm{tr}(R_{\mathrm{inv}}(S_{n}(h)))\big)=2\mathrm{tr}(R_{\mathrm{inv}}(S_{n}(h))^{2})\leq C\,\mathrm{tr}(S_{n}(h)^{6})\leq C\,\left\lVert S_{n}(h)\right\rVert_{\mathrm{op}}^{4}\,\mathrm{tr}(S_{n}(h)^{2})\to 0,$
so $Z^{\top}R_{\mathrm{inv}}(S_{n}(h))Z=o_{\mathbb{P}}(1)$.
Combining (52), (53) and (54) yields
$\ell_{n}(\theta+hr_{n}^{-1})-\ell_{n}(\theta)=h^{\top}\Xi_{n}-\frac{1}{2}h^{\top}I^{\perp}h+o_{\mathbb{P}}(1),$
as claimed.
Remark 3 (Equivalent LAN in the original $(\sigma,H)$-coordinates).
LAN is invariant under deterministic invertible linear reparametrizations of the local parameter. Let $L$ be any fixed invertible $2\times 2$ matrix and define
$\Xi_{n}^{(L)}:=L^{\top}\Xi_{n},\qquad r_{n}^{(L)}:=r_{n}L^{-1},\qquad I^{(L)}:=L^{\top}I^{\perp}L.$
Then the LAN expansion holds equivalently with $(\Xi_{n}^{(L)},I^{(L)},r_{n}^{(L)})$. In particular, choosing $L$ to map the orthogonalized coordinates back to the original parameter basis typically yields a symmetric but non-diagonal information matrix, matching standard continuous-time LAN presentations.
Remark 4.
Since $\mathbf{a}_{n}=2\ln(1/\Delta_{n})+m(H,\sigma)+o(1)$ presented in (48), in this theorem we can replace $\mathbf{a}_{n}$ by $\tilde{\mathbf{a}}_{n}:=2\ln(1/\Delta_{n})+m(H,\sigma)$ in the rate matrix $r_{n}$ (without changing the LAN expansion), because the induced change in the local perturbation $\theta_{0}+r_{n}^{-1}h$ is $o(\mathcal{T}_{n}^{-1/2})$.
6 Simulation Study
We illustrate the asymptotic behavior of the rate-matrix transformed score
$\Xi_{n}:=\frac{1}{\sqrt{\mathcal{T}_{n}}}\,M_{n}\,\nabla\ell_{n}(\theta_{0})=\binom{S_{\sigma,n}/\sqrt{\mathcal{T}_{n}}}{R_{H,n}^{\perp}/\sqrt{\mathcal{T}_{n}}},\qquad\mathcal{T}_{n}:=n\Delta_{n},$
predicted by the projected-score CLT (Proposition 5) and the LAN theorem (Theorem 2). Throughout we fix $(H,\sigma)=(0.80,1.00)$ and use the explicit integral constants $J_{0}(H,\sigma)$ and $J_{\perp}(H,\sigma)$ (Appendix A).
Projected (full-rank) limit.
Proposition 5 yields the diagonal limiting covariance
$I^{\perp}=\begin{pmatrix}\frac{\sigma^{2}}{\pi}J_{0}(H,\sigma)&0\\[2.84526pt] 0&\frac{\sigma^{4}}{4\pi}J_{\perp}(H,\sigma)\end{pmatrix}$.
For $(H,\sigma)=(0.80,1.00)$ we compute
$J_{0}=0.2819517,\qquad J_{\perp}=34.17716,\qquad\mathrm{Var}(\Xi_{\sigma})=\frac{\sigma^{2}}{\pi}J_{0}=0.0897480,\qquad\mathrm{Var}(\Xi_{H}^{\perp})=\frac{\sigma^{4}}{4\pi}J_{\perp}=2.7197322.$
The simulated joint distribution of $\Xi_{n}$ is well-approximated by the corresponding bivariate Gaussian, as shown below.
Figure 1: Projected/orthogonalized score $\Xi_{n}$ (simulation) with theoretical Gaussian contours.
Figure 2: Limiting Gaussian surface density of the projected score.
Why projection is necessary (rank-$1$ degeneracy without it).
If one only removes the explicit linear term in the $H$-score but does not project $D_{n}$ onto the orthogonal complement of $C_{n}$, the limiting covariance becomes singular (rank $1$), as explained in Appendix B. This degeneracy is visualized below.
Figure 3: Without projection: singular (rank-11) limit; the normalized pair concentrates near a line.
Appendix A Closed-form expressions for $J_{0},J_{1},J_{2}$
Recall the weight function (see (27))
$w(x):=\left(\frac{c_{H}|x|^{-p}}{1+\sigma^{2}c_{H}|x|^{-p}}\right)^{2},\qquad p=2H-1\in(1/2,1),$
and the constants (see (28)–(30))
$J_{0}(H,\sigma)$ $:=\int_{\mathbb{R}}w(x)\,dx,$
$J_{1}(H,\sigma)$ $:=\int_{\mathbb{R}}w(x)\Bigl(C_{H}-2\ln|x|\Bigr)\,dx,$
$J_{2}(H,\sigma)$ $:=\int_{\mathbb{R}}w(x)\Bigl(C_{H}-2\ln|x|\Bigr)^{2}\,dx.$
Throughout this appendix, set
$A:=\sigma^{2}c_{H}>0,\qquad L_{A}:=\ln A.$
Lemma 15.
Define, for $r$ in a neighborhood of 0,
$I(r):=\int_{0}^{\infty}\left(\frac{A\,x^{-p}}{1+A\,x^{-p}}\right)^{2}x^{r}\,dx=\int_{0}^{\infty}\frac{A^{2}x^{r-2p}}{(1+Ax^{-p})^{2}}\,dx.$
Then
$I(r)=\frac{1}{p}\,A^{(r+1)/p}\,\Gamma\!\left(\frac{r+1}{p}\right)\Gamma\!\left(2-\frac{r+1}{p}\right),$ (55)
whenever $0<\frac{r+1}{p}<2$. Moreover,
$I^{\prime}(0)=\int_{0}^{\infty}w(x)\,\ln x\,dx,\qquad I^{\prime\prime}(0)=\int_{0}^{\infty}w(x)\,(\ln x)^{2}\,dx.$
Proof.
Use the change of variables $t=Ax^{-p}$, i.e. $x=(A/t)^{1/p}$. A direct computation yields
$I(r)=\frac{1}{p}A^{(r+1)/p}\int_{0}^{\infty}\frac{t^{1-(r+1)/p}}{(1+t)^{2}}\,dt=\frac{1}{p}A^{(r+1)/p}B\!\left(2-\frac{r+1}{p},\,\frac{r+1}{p}\right),$
which gives (55) by the identity $B(x,y)=\Gamma(x)\Gamma(y)/\Gamma(x+y)$ and $\Gamma(2)=1$. Finally, differentiating under the integral sign (justified by dominated convergence for $r$ near 0) gives $I^{\prime}(0)=\int_{0}^{\infty}w(x)\ln x\,dx$ and $I^{\prime\prime}(0)=\int_{0}^{\infty}w(x)(\ln x)^{2}\,dx$. ∎
A.1 Closed form for $J_{0}$
Since $w$ is even, $J_{0}=2\int_{0}^{\infty}w(x)\,dx=2I(0)$. Lemma 15 with $r=0$ yields
$J_{0}(H,\sigma)=\frac{2}{p}\,A^{1/p}\,\Gamma\!\left(\frac{1}{p}\right)\Gamma\!\left(2-\frac{1}{p}\right)=\frac{2}{p}\,(\sigma^{2}c_{H})^{1/p}\,\Gamma\!\left(\frac{1}{p}\right)\Gamma\!\left(2-\frac{1}{p}\right).$ (56)
A.2 Closed form for $J_{1}$ and $J_{2}$
Introduce the one-sided logarithmic moments
$K_{1}:=\int_{0}^{\infty}w(x)\,\ln x\,dx=I^{\prime}(0),\qquad K_{2}:=\int_{0}^{\infty}w(x)\,(\ln x)^{2}\,dx=I^{\prime\prime}(0).$
Using evenness again,
$J_{1}=C_{H}J_{0}-4K_{1},\qquad J_{2}=C_{H}^{2}J_{0}-8C_{H}K_{1}+8K_{2}.$ (57)
To express $K_{1},K_{2}$ explicitly, differentiate $\ln I(r)$ at $r=0$. From (55),
$\ln I(r)=-\ln p+\frac{r+1}{p}\ln A+\ln\Gamma\!\left(\frac{r+1}{p}\right)+\ln\Gamma\!\left(2-\frac{r+1}{p}\right).$
Let $\psi=\Gamma^{\prime}/\Gamma$ denote the digamma function and $\psi_{1}=\psi^{\prime}$ the trigamma function. Then
$\frac{I^{\prime}(0)}{I(0)}$ $=\frac{1}{p}\Big(L_{A}+\psi(1/p)-\psi(2-1/p)\Big),$ (58)
$\frac{I^{\prime\prime}(0)}{I(0)}$ $=\Big(\frac{I^{\prime}(0)}{I(0)}\Big)^{2}+\frac{1}{p^{2}}\Big(\psi_{1}(1/p)+\psi_{1}(2-1/p)\Big).$ (59)
Since $I(0)=J_{0}/2$, we obtain
$K_{1}=\frac{J_{0}}{2}\cdot\frac{1}{p}\Big(L_{A}+\psi(1/p)-\psi(2-1/p)\Big),$ (60)
$K_{2}=\frac{J_{0}}{2}\left[\Big(\frac{I^{\prime}(0)}{I(0)}\Big)^{2}+\frac{1}{p^{2}}\Big(\psi_{1}(1/p)+\psi_{1}(2-1/p)\Big)\right].$
Combining (57) and (60) gives the closed forms:
$J_{1}(H,\sigma)=J_{0}(H,\sigma)\left[C_{H}-\frac{2}{p}\Big(L_{A}+\psi(1/p)-\psi(2-1/p)\Big)\right],$ (61)
and
$J_{2}(H,\sigma)=J_{0}(H,\sigma)\left[\Big(C_{H}-\frac{2}{p}\big(L_{A}+\psi(1/p)-\psi(2-1/p)\big)\Big)^{2}+\frac{4}{p^{2}}\big(\psi_{1}(1/p)+\psi_{1}(2-1/p)\big)\right].$ (62)
Appendix B Why two matrices $M_{n}^{(1)},M_{n}^{(2)}$: degeneracy without projection
This appendix explains why we use two successive transformations
$M_{n}^{(1)}=\begin{pmatrix}1&0\\ -\sigma\ln(\Delta_{n})&1\end{pmatrix},\qquad M_{n}^{(2)}=\begin{pmatrix}1&0\\ -\frac{\sigma a_{n}}{2}&1\end{pmatrix},\qquad M_{n}=M_{n}^{(2)}M_{n}^{(1)},$
rather than a single one, and why the projection step is necessary to obtain a full-rank ($2\times 2$) limiting information.
B.1 Step 1: remove the explicit linear term in the $H$-score
Recall the exact decomposition (see Lemma 1)
$S_{H,n}=\sigma\ln(\Delta_{n})\,S_{\sigma,n}+R_{H,n},\qquad R_{H,n}:=\frac{\gamma_{n}}{2}\Big(Z^{\top}D_{n}Z-\mathrm{tr}(D_{n})\Big).$ (63)
Therefore,
$M_{n}^{(1)}\binom{S_{\sigma,n}}{S_{H,n}}=\binom{S_{\sigma,n}}{R_{H,n}}$.
This first step produces a pair of centered Gaussian quadratic forms.
B.2 Step 2: projection and orthogonalization
Define the projection coefficient and orthogonalized matrix (Definition 2)
$a_{n}:=\frac{\mathrm{tr}(C_{n}D_{n})}{\mathrm{tr}(C_{n}^{2})},\qquad D_{n}^{\perp}:=D_{n}-a_{n}C_{n},\qquad\mathrm{tr}(C_{n}D_{n}^{\perp})=0,$
and the orthogonalized remainder
$R_{H,n}^{\perp}:=\frac{\gamma_{n}}{2}\Big(Z^{\top}D_{n}^{\perp}Z-\mathrm{tr}(D_{n}^{\perp})\Big)=R_{H,n}-\frac{\sigma a_{n}}{2}\,S_{\sigma,n}.$
This corresponds exactly to the second transformation
$M_{n}^{(2)}\binom{S_{\sigma,n}}{R_{H,n}}=\binom{S_{\sigma,n}}{R_{H,n}^{\perp}}$.
By Wick’s identity for Gaussian quadratic forms and $\mathrm{tr}(C_{n}D_{n}^{\perp})=0$, one has the exact orthogonality
$\mathrm{Cov}(S_{\sigma,n},R_{H,n}^{\perp})=0\qquad\text{for every }n$
(cf. Lemma 9).
B.3 Singular covariance without the projection step
Let $\mathcal{T}_{n}:=n\Delta_{n}$ and $L_{n}:=\ln(1/\Delta_{n})=|\ln\Delta_{n}|$. Consider the normalized pair after only the first transformation:
$U_{n}:=\binom{\displaystyle\frac{S_{\sigma,n}}{\sqrt{\mathcal{T}_{n}}}}{\displaystyle\frac{R_{H,n}}{L_{n}\sqrt{\mathcal{T}_{n}}}}$.
Proposition 6.
Recall the trace asymptotics proved in Lemma 5:
$\mathrm{tr}(C_{n}^{2})$ $=\frac{n}{2\pi}\Delta_{n}^{1-2p}\big(J_{0}+o(1)\big),$
$\mathrm{tr}(C_{n}D_{n})$ $=\frac{n}{2\pi}\Delta_{n}^{1-2p}\big(2L_{n}J_{0}+J_{1}+o(1)\big),$
$\mathrm{tr}(D_{n}^{2})$ $=\frac{n}{2\pi}\Delta_{n}^{1-2p}\big(4L_{n}^{2}J_{0}+4L_{n}J_{1}+J_{2}+o(1)\big).$
Combining these with the operator–Frobenius negligibility (which justifies the quadratic-form CLT), the vector $U_{n}$ converges jointly to a centered Gaussian vector with covariance matrix
$\Sigma^{(1)}=\frac{J_{0}(H,\sigma)}{\pi}\begin{pmatrix}\sigma^{2}&\sigma^{3}\\ \sigma^{3}&\sigma^{4}\end{pmatrix},\qquad\mathrm{rank}(\Sigma^{(1)})=1.$
Equivalently, the two components are asymptotically perfectly correlated:
$\mathrm{Corr}\!\left(\frac{S_{\sigma,n}}{\sqrt{\mathcal{T}_{n}}},\frac{R_{H,n}}{L_{n}\sqrt{\mathcal{T}_{n}}}\right)\longrightarrow 1.$
Proof.
Write $Q(A):=Z^{\top}AZ-\mathrm{tr}(A)$. Then
$S_{\sigma,n}=\frac{\gamma_{n}}{\sigma}Q(C_{n}),\qquad R_{H,n}=\frac{\gamma_{n}}{2}Q(D_{n}),\qquad\gamma_{n}=\sigma^{2}\Delta_{n}^{p}.$
By Wick’s identity,
$\mathrm{Cov}(Q(A),Q(B))=2\mathrm{tr}(AB),\qquad\mathrm{Var}(Q(A))=2\mathrm{tr}(A^{2}).$
Hence
$\mathrm{Var}\!\left(\frac{S_{\sigma,n}}{\sqrt{\mathcal{T}_{n}}}\right)=\frac{\gamma_{n}^{2}}{\sigma^{2}\mathcal{T}_{n}}\cdot 2\mathrm{tr}(C_{n}^{2})\to\frac{\sigma^{2}}{\pi}J_{0},$
and similarly
$\mathrm{Var}\!\left(\frac{R_{H,n}}{L_{n}\sqrt{\mathcal{T}_{n}}}\right)=\frac{\gamma_{n}^{2}}{4L_{n}^{2}\mathcal{T}_{n}}\cdot 2\mathrm{tr}(D_{n}^{2})\to\frac{\sigma^{4}}{\pi}J_{0}.$
For the covariance,
$\mathrm{Cov}\!\left(\frac{S_{\sigma,n}}{\sqrt{\mathcal{T}_{n}}},\frac{R_{H,n}}{L_{n}\sqrt{\mathcal{T}_{n}}}\right)=\frac{\gamma_{n}^{2}}{2\sigma L_{n}\mathcal{T}_{n}}\cdot 2\mathrm{tr}(C_{n}D_{n})=\frac{\gamma_{n}^{2}}{\sigma L_{n}\mathcal{T}_{n}}\mathrm{tr}(C_{n}D_{n}).$
Using $\gamma_{n}^{2}=\sigma^{4}\Delta_{n}^{2p}$ and $\mathcal{T}_{n}=n\Delta_{n}$ together with the stated asymptotic for $\mathrm{tr}(C_{n}D_{n})$ yields
$\mathrm{Cov}\!\left(\frac{S_{\sigma,n}}{\sqrt{\mathcal{T}_{n}}},\frac{R_{H,n}}{L_{n}\sqrt{\mathcal{T}_{n}}}\right)=\frac{\sigma^{4}\Delta_{n}^{2p}}{\sigma L_{n}\,n\Delta_{n}}\cdot\frac{n}{2\pi}\Delta_{n}^{1-2p}\big(2L_{n}J_{0}+o(L_{n})\big)\to\frac{\sigma^{3}}{\pi}J_{0}.$
Thus the limiting covariance matrix is exactly $\Sigma^{(1)}$, whose determinant is 0, hence rank $1$. ∎
Appendix C The regime $1/2<H<3/4$
This appendix treats the subcritical long-memory regime $1/2<H<3/4$, i.e.
$p:=2H-1\in(0,1/2),\qquad\Delta_{n}=n^{-\alpha}\ \ (\alpha\in(0,1)),\qquad\gamma_{n}=\sigma^{2}\Delta_{n}^{p}\to 0.$
In contrast to the supercritical case $H>3/4$ (where $f_{H}^{2}\notin L^{1}$), here we have
$f_{H}\in L^{2}([-\pi,\pi]),\qquad\dot{f}_{H}\in L^{2}([-\pi,\pi]),$
and thus the relevant Toeplitz trace functionals are of order $n$. Here we only deal with the CLT for the score function with rate matrix but not the LAN property which can be obtained easily as in the section 5.
C.1 Trace approximations at order $n$.
Recall $A_{n}:=I_{n}+\gamma_{n}T_{n}(H)=T_{n}(a_{n})$ with $a_{n}(\lambda):=1+\gamma_{n}f_{H}(\lambda)$ and
$C_{n}:=A_{n}^{-1/2}T_{n}(H)A_{n}^{-1/2},\qquad D_{n}:=A_{n}^{-1/2}\dot{T}_{n}(H)A_{n}^{-1/2}.$
Define the (triangular-array) symbols
$g_{n}(\lambda):=\frac{f_{H}(\lambda)}{a_{n}(\lambda)},\qquad h_{n}(\lambda):=\frac{\dot{f}_{H}(\lambda)}{a_{n}(\lambda)}.$
Lemma 16.
Assume $1/2<H<3/4$ (equivalently $p\in(0,1/2)$) and $\gamma_{n}=\sigma^{2}\Delta_{n}^{p}\to 0$. Then
$\mathrm{tr}(C_{n}^{2})$ $=\frac{n}{2\pi}\int_{-\pi}^{\pi}g_{n}(\lambda)^{2}\,d\lambda+o(n),$ (64)
$\mathrm{tr}(C_{n}D_{n})$ $=\frac{n}{2\pi}\int_{-\pi}^{\pi}g_{n}(\lambda)h_{n}(\lambda)\,d\lambda+o(n),$ (65)
$\mathrm{tr}(D_{n}^{2})$ $=\frac{n}{2\pi}\int_{-\pi}^{\pi}h_{n}(\lambda)^{2}\,d\lambda+o(n).$ (66)
Proof.
The key point is that for $p\in(0,1/2)$ we have $f_{H},\dot{f}_{H}\in L^{2}([-\pi,\pi])$, hence $g_{n},h_{n}\in L^{2}$ uniformly because $a_{n}(\lambda)\geq 1$. In this $L^{2}$-setting, classical Szegő/Avram-type trace theorems for products of Toeplitz matrices yield the order-$n$ trace approximation with $o(n)$ remainder, see e.g. Avram [1] and standard Toeplitz references such as Grenander–Szegő [5] or Böttcher–Silbermann [2]. Applying these results to the sandwiched forms corresponding to $C_{n},D_{n}$ (equivalently, to the symbols $g_{n},h_{n}$) gives (64)–(66). ∎
Lemma 17.
Under $1/2<H<3/4$ (so $p\in(0,1/2)$) and $\gamma_{n}\to 0$,
$\int_{-\pi}^{\pi}g_{n}(\lambda)^{2}\,d\lambda$ $\longrightarrow\int_{-\pi}^{\pi}f_{H}(\lambda)^{2}\,d\lambda,$ (67)
$\int_{-\pi}^{\pi}g_{n}(\lambda)h_{n}(\lambda)\,d\lambda$ $\longrightarrow\int_{-\pi}^{\pi}f_{H}(\lambda)\dot{f}_{H}(\lambda)\,d\lambda,$ (68)
$\int_{-\pi}^{\pi}h_{n}(\lambda)^{2}\,d\lambda$ $\longrightarrow\int_{-\pi}^{\pi}\dot{f}_{H}(\lambda)^{2}\,d\lambda.$ (69)
Proof.
Since $a_{n}(\lambda)=1+\gamma_{n}f_{H}(\lambda)\to 1$ pointwise and $a_{n}(\lambda)\geq 1$, we have $|g_{n}|\leq|f_{H}|$ and $|h_{n}|\leq|\dot{f}_{H}|$. Because $f_{H},\dot{f}_{H}\in L^{2}([-\pi,\pi])$ for $p\in(0,1/2)$, dominated convergence yields (67)–(69). ∎
C.2 Operator/Frobenius ratios in the subcritical regime
The op/F condition (12) required by Lemma 2 follows from the fact that the operator norm of long-memory Toeplitz matrices grows like $n^{p}$ (up to log-factors for $\dot{f}_{H}$), while the Frobenius norms scale like $\sqrt{n}$.
Lemma 18.
Fix $1/2<H<3/4$ and set $p:=2H-1\in(0,1/2)$. Then
$\frac{\left\lVert C_{n}\right\rVert_{\mathrm{op}}}{\left\lVert C_{n}\right\rVert_{F}}\longrightarrow 0,\qquad\frac{\left\lVert D_{n}\right\rVert_{\mathrm{op}}}{\left\lVert D_{n}\right\rVert_{F}}\longrightarrow 0.$
Proof.
Since $A_{n}\succeq I_{n}$, one has $\left\lVert A_{n}^{-1/2}\right\rVert_{\mathrm{op}}\leq 1$, hence
$\left\lVert C_{n}\right\rVert_{\mathrm{op}}\leq\left\lVert T_{n}(H)\right\rVert_{\mathrm{op}},\qquad\left\lVert D_{n}\right\rVert_{\mathrm{op}}\leq\left\lVert\dot{T}_{n}(H)\right\rVert_{\mathrm{op}}.$
For long-memory Toeplitz matrices with Fisher–Hartwig exponent $p\in(0,1)$, it is standard that
$\left\lVert T_{n}(H)\right\rVert_{\mathrm{op}}=O(n^{p}),\qquad\left\lVert\dot{T}_{n}(H)\right\rVert_{\mathrm{op}}=O(n^{p}\log n),$
see e.g. Böttcher–Silbermann [2] or Gray [4] (the $\log n$ comes from the logarithmic factor in $\dot{f}_{H}$). On the other hand, Lemma 16 and Lemma 17 give
$\left\lVert C_{n}\right\rVert_{F}^{2}=\mathrm{tr}(C_{n}^{2})=\frac{n}{2\pi}\int_{-\pi}^{\pi}f_{H}(\lambda)^{2}\,d\lambda+o(n)\asymp n,$
and similarly $\left\lVert D_{n}\right\rVert_{F}^{2}=\mathrm{tr}(D_{n}^{2})\asymp n$. Therefore
$\frac{\left\lVert C_{n}\right\rVert_{\mathrm{op}}}{\left\lVert C_{n}\right\rVert_{F}}\lesssim\frac{n^{p}}{\sqrt{n}}=n^{p-\frac{1}{2}}\to 0,\qquad\frac{\left\lVert D_{n}\right\rVert_{\mathrm{op}}}{\left\lVert D_{n}\right\rVert_{F}}\lesssim\frac{n^{p}\log n}{\sqrt{n}}=n^{p-\frac{1}{2}}\log n\to 0,$
because $p<1/2$. ∎
C.3 Joint CLT for the transformed score vector and the information matrix
Recall from Lemma 1 the exact quadratic-form identities (9) and (10)–(11):
$S_{\sigma,n}=\frac{\gamma_{n}}{\sigma}\Big(Z^{\top}C_{n}Z-\mathrm{tr}(C_{n})\Big),\qquad S_{H,n}=\sigma\ln(\Delta_{n})\,S_{\sigma,n}+R_{H,n},\quad R_{H,n}=\frac{\gamma_{n}}{2}\Big(Z^{\top}D_{n}Z-\mathrm{tr}(D_{n})\Big).$
Let the subcritical normalization be
$v_{n}:=\sqrt{n}\,\Delta_{n}^{p}.$
Proposition 7.
Assume $1/2<H<3/4$ and $\Delta_{n}=n^{-\alpha}$ with $\alpha\in(0,1)$, so that $\gamma_{n}=\sigma^{2}\Delta_{n}^{p}\to 0$. Then
$\frac{1}{v_{n}}\binom{S_{\sigma,n}}{R_{H,n}}\ \Rightarrow\ \mathcal{N}\!\left(0,\ I^{(<)}(\sigma,H)\right),$
where
$I^{(<)}(\sigma,H):=\begin{pmatrix}\displaystyle\frac{2\sigma^{2}}{\pi}\int_{0}^{\pi}f_{H}(\lambda)^{2}\,d\lambda&\displaystyle\frac{\sigma^{3}}{\pi}\int_{0}^{\pi}f_{H}(\lambda)\dot{f}_{H}(\lambda)\,d\lambda\\[8.53581pt] \displaystyle\frac{\sigma^{3}}{\pi}\int_{0}^{\pi}f_{H}(\lambda)\dot{f}_{H}(\lambda)\,d\lambda&\displaystyle\frac{\sigma^{4}}{2\pi}\int_{0}^{\pi}\dot{f}_{H}(\lambda)^{2}\,d\lambda\end{pmatrix}$.
Equivalently, since $M_{n}^{(1)}$ is defined in (50) and $M_{n}^{(1)}(S_{\sigma,n},S_{H,n})^{\top}=(S_{\sigma,n},R_{H,n})^{\top}$ by (10), one has
$\frac{1}{v_{n}}\,M_{n}^{(1)}\binom{S_{\sigma,n}}{S_{H,n}}\ \Rightarrow\ \mathcal{N}\!\left(0,\ I^{(<)}(\sigma,H)\right).$
Moreover, $I^{(<)}(\sigma,H)$ is non-singular in this regime, hence no second projection is needed.
Proof.
First we construct the one-dimensional CLT via op/F structure. Fix $u=(u_{1},u_{2})\in\mathbb{R}^{2}$ and set
$M_{n,u}:=u_{1}\sigma^{-1}C_{n}+u_{2}\tfrac{1}{2}D_{n},\qquad Q_{n}(M_{n,u}):=Z^{\top}M_{n,u}Z-\mathrm{tr}(M_{n,u}).$
Then
$u_{1}\frac{S_{\sigma,n}}{v_{n}}+u_{2}\frac{R_{H,n}}{v_{n}}=\frac{\gamma_{n}}{v_{n}}\,Q_{n}(M_{n,u}).$
By Lemma 18, $\left\lVert M_{n,u}\right\rVert_{\mathrm{op}}/\left\lVert M_{n,u}\right\rVert_{F}\to 0$, so Lemma 2 yields
$\frac{Q_{n}(M_{n,u})}{\sqrt{2}\,\left\lVert M_{n,u}\right\rVert_{F}}\Rightarrow\mathcal{N}(0,1).$
Now we should identify the asymptotic variance. By definition of $M_{u,v}$,
$\mathrm{Var}\!\left(\frac{\gamma_{n}}{v_{n}}Q_{n}(M_{n,u})\right)=\frac{\gamma_{n}^{2}}{v_{n}^{2}}\cdot 2\,\left\lVert M_{n,u}\right\rVert_{F}^{2}=\frac{\gamma_{n}^{2}}{n\Delta_{n}^{2p}}\cdot 2\,\mathrm{tr}(M_{n,u}^{2}).$
Using bilinearity and Wick’s identity, $\mathrm{tr}(M_{n,u}^{2})=u_{1}^{2}\sigma^{-2}\mathrm{tr}(C_{n}^{2})+u_{1}u_{2}\sigma^{-1}\mathrm{tr}(C_{n}D_{n})+\frac{u_{2}^{2}}{4}\mathrm{tr}(D_{n}^{2})$. Lemma 16 and Lemma 17 give
$\frac{1}{n}\mathrm{tr}(C_{n}^{2})\to\frac{1}{2\pi}\int_{-\pi}^{\pi}f_{H}(\lambda)^{2}\,d\lambda,\quad\frac{1}{n}\mathrm{tr}(C_{n}D_{n})\to\frac{1}{2\pi}\int_{-\pi}^{\pi}f_{H}(\lambda)\dot{f}_{H}(\lambda)\,d\lambda,$
$\frac{1}{n}\mathrm{tr}(D_{n}^{2})\to\frac{1}{2\pi}\int_{-\pi}^{\pi}\dot{f}_{H}(\lambda)^{2}\,d\lambda.$
Finally, since $\gamma_{n}^{2}/(n\Delta_{n}^{2p})=\sigma^{4}\Delta_{n}^{2p}/(n\Delta_{n}^{2p})=\sigma^{4}/n$, the prefactor combines with the $n$-scale of the traces to yield the stated limits, and rewriting $\pi,\pi]$ as $2[0,\pi]$ gives exactly the covariance matrix $I^{(<)}(\sigma,H)$.
At last we will construct the joint convergence using the Cramér–Wold theorem: every fixed linear combination converges to a centered normal with the variance prescribed by $I^{(<)}(\sigma,H)$, hence the vector converges jointly. ∎
Appendix D The regime $0<H<\tfrac{1}{2}$
In the fBm-dominated regime $0<H<\frac{1}{2}$, we also consider the important part–the CLT of the score function. When this one has been established, we can prove the LAN property the same way as in [9]. In this case it is more convenient to work with the small parameter
$\varepsilon_{n}:=\Delta_{n}^{\,1-2H}\ \downarrow\ 0,$
rather than $\gamma_{n}=\sigma^{2}\Delta_{n}^{2H-1}\to\infty$. We start from the exact covariance of the observed increment vector $X_{n}=(X_{1,n},\dots,X_{n,n})^{\top}$:
$V_{n}(\sigma,H)=\mathrm{Var}(X_{n})=\sigma^{2}\Delta_{n}^{2H}\,T_{n}(H)+\Delta_{n}I_{n},$ (70)
where $T_{n}(H)=T_{n}(f_{H})$ is the Toeplitz covariance matrix of the standard fGn increment sequence.
Factoring out $\Delta_{n}^{2H}$ yields the equivalent representation
$V_{n}(\sigma,H)=\Delta_{n}^{2H}\,B_{n}(\sigma,H),\qquad B_{n}(\sigma,H):=\sigma^{2}T_{n}(H)+\varepsilon_{n}I_{n}.$ (71)
Note that
$\gamma_{n}=\sigma^{2}\Delta_{n}^{2H-1}=\frac{\sigma^{2}}{\varepsilon_{n}},\qquad A_{n}:=I_{n}+\gamma_{n}T_{n}(H)=\varepsilon_{n}^{-1}B_{n}(\sigma,H),$
so the two normalizations $V_{n}=\Delta_{n}A_{n}$ and $V_{n}=\Delta_{n}^{2H}B_{n}$ are identical.
Under $\mathbb{P}_{\sigma,H}$ define
$Z:=V_{n}(\sigma,H)^{-1/2}X_{n}=\Delta_{n}^{-H}\,B_{n}(\sigma,H)^{-1/2}X_{n}\sim\mathcal{N}(0,I_{n}).$ (72)
This is the same Gaussian vector $Z$ as in Lemma 1; we only rewrote the whitening in terms of $B_{n}$. Now it is natural to define
$\widetilde{C}_{n}:=B_{n}(\sigma,H)^{-1/2}\,T_{n}(H)\,B_{n}(\sigma,H)^{-1/2},\qquad\widetilde{D}_{n}:=B_{n}(\sigma,H)^{-1/2}\,\dot{T}_{n}(H)\,B_{n}(\sigma,H)^{-1/2}.$ (73)
These matrices are related to the matrices $C_{n},D_{n}$ in Lemma 1 by
$C_{n}=\varepsilon_{n}\,\widetilde{C}_{n},\qquad D_{n}=\varepsilon_{n}\,\widetilde{D}_{n},$ (74)
since $A_{n}^{-1/2}=\sqrt{\varepsilon_{n}}\,B_{n}^{-1/2}$. We will write the real formula of the score function:
Lemma 19.
Assume $0<H<\frac{1}{2}$. With $Z$ as in (72) and $(\widetilde{C}_{n},\widetilde{D}_{n})$ as in (73), the score representations (9)–(11) can be rewritten as
$S_{\sigma,n}$ $=\sigma\Big(Z^{\top}\widetilde{C}_{n}Z-\mathrm{tr}(\widetilde{C}_{n})\Big),$ (75)
$R_{H,n}$ $=\frac{\sigma^{2}}{2}\Big(Z^{\top}\widetilde{D}_{n}Z-\mathrm{tr}(\widetilde{D}_{n})\Big),$ (76)
$S_{H,n}$ $=\sigma\ln(\Delta_{n})\,S_{\sigma,n}+R_{H,n},$ (77)
where $R_{H,n}$ is the same remainder as in (11).
Proof.
Using (74) and $\gamma_{n}=\sigma^{2}/\varepsilon_{n}$, the identity (9) becomes
$S_{\sigma,n}=\frac{\gamma_{n}}{\sigma}\Big(Z^{\top}(\varepsilon_{n}\widetilde{C}_{n})Z-\mathrm{tr}(\varepsilon_{n}\widetilde{C}_{n})\Big)=\sigma\Big(Z^{\top}\widetilde{C}_{n}Z-\mathrm{tr}(\widetilde{C}_{n})\Big),$
which is (75). The same argument applied to (11) gives (76). Finally, (77) is exactly (10). ∎
D.1 Toeplitz trace approximations for $(\widetilde{C}_{n},\widetilde{D}_{n})$
Define the (triangular-array) symbols
$\widetilde{c}_{n}(\lambda):=\frac{f_{H}(\lambda)}{\sigma^{2}f_{H}(\lambda)+\varepsilon_{n}},\qquad\widetilde{d}_{n}(\lambda):=\frac{\dot{f}_{H}(\lambda)}{\sigma^{2}f_{H}(\lambda)+\varepsilon_{n}},\qquad\lambda\in[-\pi,\pi].$ (78)
Then $0\leq\widetilde{c}_{n}(\lambda)\leq\sigma^{-2}$ and $\widetilde{d}_{n}(\lambda)=\widetilde{c}_{n}(\lambda)\,b_{H}(\lambda)$ with
$b_{H}(\lambda):=\partial_{H}\log f_{H}(\lambda)=\frac{\dot{f}_{H}(\lambda)}{f_{H}(\lambda)}.$ (79)
Lemma 20.
Assume $0<H<\frac{1}{2}$ and Lemma 3. Then
$\mathrm{tr}(\widetilde{C}_{n}^{2})$ $=\frac{n}{2\pi}\int_{-\pi}^{\pi}\widetilde{c}_{n}(\lambda)^{2}\,d\lambda+o(n),$ (80)
$\mathrm{tr}(\widetilde{C}_{n}\widetilde{D}_{n})$ $=\frac{n}{2\pi}\int_{-\pi}^{\pi}\widetilde{c}_{n}(\lambda)\widetilde{d}_{n}(\lambda)\,d\lambda+o(n),$ (81)
$\mathrm{tr}(\widetilde{D}_{n}^{2})$ $=\frac{n}{2\pi}\int_{-\pi}^{\pi}\widetilde{d}_{n}(\lambda)^{2}\,d\lambda+o(n).$ (82)
Proof.
Since $H<1/2$, the spectral density $f_{H}$ is bounded on $\pi,\pi]$ and strictly positive on $\pi,\pi]\setminus\{0\}$. Moreover, by Lemma 3, $b_{H}(\lambda)=\partial_{H}\log f_{H}(\lambda)$ has at most a logarithmic singularity at $\lambda=0$, hence $b_{H}\in L^{2}([-\pi,\pi])$. Therefore $\widetilde{c}_{n}$ is uniformly bounded and $|\widetilde{d}_{n}(\lambda)|\leq\sigma^{-2}|b_{H}(\lambda)|$ is square-integrable uniformly in $n$.
In this $L^{2}$ setting, standard Szegő/Avram-type trace theorems for products of Toeplitz matrices apply to the triangular-array symbols $\widetilde{c}_{n},\widetilde{d}_{n}$ and yield (80)–(82) with $o(n)$ remainder. ∎
D.2 Limits of the trace integrals and the constants $T_{1}(H),T_{2}(H)$
Define
$T_{1}(H):=\frac{1}{2\pi}\int_{-\pi}^{\pi}b_{H}(\lambda)\,d\lambda,\qquad T_{2}(H):=\frac{1}{4\pi}\int_{-\pi}^{\pi}b_{H}(\lambda)^{2}\,d\lambda.$ (83)
Lemma 21 (Integral limits).
Assume $0<H<\frac{1}{2}$ and $\varepsilon_{n}\to 0$. Then
$\int_{-\pi}^{\pi}\widetilde{c}_{n}(\lambda)^{2}\,d\lambda$ $\longrightarrow\frac{2\pi}{\sigma^{4}},$ (84)
$\int_{-\pi}^{\pi}\widetilde{c}_{n}(\lambda)\widetilde{d}_{n}(\lambda)\,d\lambda$ $\longrightarrow\frac{2\pi}{\sigma^{4}}\,T_{1}(H),$ (85)
$\int_{-\pi}^{\pi}\widetilde{d}_{n}(\lambda)^{2}\,d\lambda$ $\longrightarrow\frac{4\pi}{\sigma^{4}}\,T_{2}(H).$ (86)
Proof.
For each $\lambda\neq 0$, since $f_{H}(\lambda)>0$ we have
$\widetilde{c}_{n}(\lambda)=\frac{f_{H}(\lambda)}{\sigma^{2}f_{H}(\lambda)+\varepsilon_{n}}\ \longrightarrow\ \frac{1}{\sigma^{2}}.$
Moreover $0\leq\widetilde{c}_{n}\leq\sigma^{-2}$, hence dominated convergence gives (84).
Next, $\widetilde{d}_{n}(\lambda)=\widetilde{c}_{n}(\lambda)\,b_{H}(\lambda)$ and $b_{H}\in L^{1}([-\pi,\pi])$ (log singularity), so $|\widetilde{c}_{n}b_{H}|\leq\sigma^{-2}|b_{H}|$ is integrable and dominated convergence yields (85). Finally, $b_{H}^{2}\in L^{1}([-\pi,\pi])$ and $|\widetilde{d}_{n}|^{2}=\widetilde{c}_{n}^{2}b_{H}^{2}\leq\sigma^{-4}b_{H}^{2}$, hence dominated convergence yields (86). ∎
D.3 Quotien of Two Norms
With the trace approximation, we can easily construct the op/F
Lemma 22.
Fix $0<H<\frac{1}{2}$. Then
$\frac{\left\lVert\widetilde{C}_{n}\right\rVert_{\mathrm{op}}}{\left\lVert\widetilde{C}_{n}\right\rVert_{F}}\longrightarrow 0,\qquad\frac{\left\lVert\widetilde{D}_{n}\right\rVert_{\mathrm{op}}}{\left\lVert\widetilde{D}_{n}\right\rVert_{F}}\longrightarrow 0.$
Proof.
First we will calculate the operator norm bounds. Since $B_{n}=\sigma^{2}T_{n}(H)+\varepsilon_{n}I_{n}$ is a polynomial in $T_{n}(H)$, the matrices $B_{n}$ and $T_{n}(H)$ commute and are simultaneously diagonalizable. Hence the eigenvalues of $\widetilde{C}_{n}=B_{n}^{-1/2}T_{n}(H)B_{n}^{-1/2}$ are of the form $\lambda/(\sigma^{2}\lambda+\varepsilon_{n})$ with $\lambda\geq 0$, so
$\left\lVert\widetilde{C}_{n}\right\rVert_{\mathrm{op}}\leq\sup_{\lambda\geq 0}\frac{\lambda}{\sigma^{2}\lambda+\varepsilon_{n}}\leq\frac{1}{\sigma^{2}}.$
For $\widetilde{D}_{n}=B_{n}^{-1/2}\dot{T}_{n}(H)B_{n}^{-1/2}$, set
$G_{n}:=T_{n}(H)^{1/2}B_{n}^{-1/2}=B_{n}^{-1/2}T_{n}(H)^{1/2},\qquad E_{n}:=T_{n}(H)^{-1/2}\dot{T}_{n}(H)\,T_{n}(H)^{-1/2}.$
Then $\widetilde{D}_{n}=G_{n}E_{n}G_{n}$ and thus
$\left\lVert\widetilde{D}_{n}\right\rVert_{\mathrm{op}}\leq\left\lVert G_{n}\right\rVert_{\mathrm{op}}^{2}\,\left\lVert E_{n}\right\rVert_{\mathrm{op}}=\left\lVert\widetilde{C}_{n}\right\rVert_{\mathrm{op}}\,\left\lVert E_{n}\right\rVert_{\mathrm{op}}\leq\frac{1}{\sigma^{2}}\,\left\lVert E_{n}\right\rVert_{\mathrm{op}}.$
By the same argument as in Lemma 7 (Toeplitz log-derivative symbol and Toeplitz–Hankel sandwich), one has $\left\lVert E_{n}\right\rVert_{\mathrm{op}}=O(\log n)$, hence $\left\lVert\widetilde{D}_{n}\right\rVert_{\mathrm{op}}=O(\log n)$.
Now we will study the Frobenius norms. By Lemma 20 and Lemma 21,
$\left\lVert\widetilde{C}_{n}\right\rVert_{F}^{2}=\mathrm{tr}(\widetilde{C}_{n}^{2})=\frac{n}{2\pi}\int_{-\pi}^{\pi}\widetilde{c}_{n}(\lambda)^{2}\,d\lambda+o(n)=\frac{n}{\sigma^{4}}\bigl(1+o(1)\bigr),$
so $\left\lVert\widetilde{C}_{n}\right\rVert_{F}\asymp\sqrt{n}$. Similarly, (82) and (86) give $\left\lVert\widetilde{D}_{n}\right\rVert_{F}\asymp\sqrt{n}$.
When the two norm are well prepared, we can easily obtain
$\frac{\left\lVert\widetilde{C}_{n}\right\rVert_{\mathrm{op}}}{\left\lVert\widetilde{C}_{n}\right\rVert_{F}}\lesssim\frac{1}{\sqrt{n}}\to 0,\qquad\frac{\left\lVert\widetilde{D}_{n}\right\rVert_{\mathrm{op}}}{\left\lVert\widetilde{D}_{n}\right\rVert_{F}}\lesssim\frac{\log n}{\sqrt{n}}\to 0.$
D.4 Joint CLT the same as pure-fGn information matrix
Define the linear transformation removing the deterministic linear term in $S_{H,n}$:
$M_{n}^{(1)}:=\begin{pmatrix}1&0\\ -\sigma\ln(\Delta_{n})&1\end{pmatrix},\qquad M_{n}^{(1)}\binom{S_{\sigma,n}}{S_{H,n}}=\binom{S_{\sigma,n}}{R_{H,n}}$, (87)
cf. Lemma 1 and (77).
Proposition 8.
Assume $0<H<\frac{1}{2}$ and Lemma 3. Then, as $n\to\infty$,
$\frac{1}{\sqrt{n}}\,M_{n}^{(1)}\binom{S_{\sigma,n}}{S_{H,n}}=\binom{\frac{S_{\sigma,n}}{\sqrt{n}}}{\frac{R_{H,n}}{\sqrt{n}}}\ \Rightarrow\ \mathcal{N}\!\left(0,\ I_{\mathrm{pure}}(\sigma,H)\right),$
where
$I_{\mathrm{pure}}(\sigma,H)=\begin{pmatrix}\frac{2}{\sigma^{2}}&\frac{1}{\sigma}\,T_{1}(H)\\[2.84526pt] \frac{1}{\sigma}\,T_{1}(H)&T_{2}(H)\end{pmatrix}$, (88)
with $T_{1}(H),T_{2}(H)$ defined in (83).
Proof.
By Lemma 19,
$S_{\sigma,n}=\sigma\Big(Z^{\top}\widetilde{C}_{n}Z-\mathrm{tr}(\widetilde{C}_{n})\Big),\qquad R_{H,n}=\frac{\sigma^{2}}{2}\Big(Z^{\top}\widetilde{D}_{n}Z-\mathrm{tr}(\widetilde{D}_{n})\Big),\qquad Z\sim\mathcal{N}(0,I_{n}).$
Using $\mathrm{Var}(Z^{\top}AZ-\mathrm{tr}A)=2\mathrm{tr}(A^{2})$ and Lemmas 2021, we obtain
$\mathrm{Var}\!\Big(\frac{S_{\sigma,n}}{\sqrt{n}}\Big)=\frac{\sigma^{2}}{n}\cdot 2\mathrm{tr}(\widetilde{C}_{n}^{2})=\frac{2\sigma^{2}}{n}\left(\frac{n}{2\pi}\int\widetilde{c}_{n}^{2}+o(n)\right)\to\frac{2\sigma^{2}}{2\pi}\cdot\frac{2\pi}{\sigma^{4}}=\frac{2}{\sigma^{2}},$
and
$\mathrm{Var}\!\Big(\frac{R_{H,n}}{\sqrt{n}}\Big)=\frac{\sigma^{4}}{4n}\cdot 2\mathrm{tr}(\widetilde{D}_{n}^{2})=\frac{\sigma^{4}}{2n}\left(\frac{n}{2\pi}\int\widetilde{d}_{n}^{2}+o(n)\right)\to\frac{\sigma^{4}}{2\cdot 2\pi}\cdot\frac{4\pi}{\sigma^{4}}=T_{2}(H),$
where the last equality uses (86) and (83).
By Wick’s identity,
$\mathrm{Cov}\!\Big(\frac{S_{\sigma,n}}{\sqrt{n}},\frac{R_{H,n}}{\sqrt{n}}\Big)=\frac{\sigma\cdot(\sigma^{2}/2)}{n}\cdot 2\,\mathrm{tr}(\widetilde{C}_{n}\widetilde{D}_{n})=\frac{\sigma^{3}}{n}\mathrm{tr}(\widetilde{C}_{n}\widetilde{D}_{n}).$
Using (81) and (85),
$\frac{1}{n}\mathrm{tr}(\widetilde{C}_{n}\widetilde{D}_{n})=\frac{1}{2\pi}\int\widetilde{c}_{n}\widetilde{d}_{n}+o(1)\to\frac{1}{2\pi}\cdot\frac{2\pi}{\sigma^{4}}\,T_{1}(H)=\frac{T_{1}(H)}{\sigma^{4}},$
hence the covariance limit equals $\sigma^{3}\cdot\sigma^{-4}T_{1}(H)=\sigma^{-1}T_{1}(H)$.
For any fixed $u=(u_{1},u_{2})\in\mathbb{R}^{2}$,
$u_{1}\frac{S_{\sigma,n}}{\sqrt{n}}+u_{2}\frac{R_{H,n}}{\sqrt{n}}=\frac{1}{\sqrt{n}}\Big(Z^{\top}\widetilde{M}_{n,u}Z-\mathrm{tr}(\widetilde{M}_{n,u})\Big),\qquad\widetilde{M}_{n,u}:=u_{1}\sigma\,\widetilde{C}_{n}+u_{2}\frac{\sigma^{2}}{2}\,\widetilde{D}_{n}.$
By Lemma 22, we have $\left\lVert\widetilde{M}_{n,u}\right\rVert_{\mathrm{op}}/\left\lVert\widetilde{M}_{n,u}\right\rVert_{F}\to 0$, so Lemma 2 yields a one-dimensional normal limit for every $u$. Therefore the vector converges jointly to a centered Gaussian with covariance matrix (88). ∎
Remark 5 (Consistency with the pure fGn experiment).
The limiting covariance (88) coincides with the Fisher information of the pure fGn model in the regime $H<1/2$. The Brownian component enters only through the vanishing regularization $\varepsilon_{n}I_{n}$ in $B_{n}=\sigma^{2}T_{n}(H)+\varepsilon_{n}I_{n}$.